### v1

# Stick to your Role! Stability of Personal Values Expressed in Large Language Models

Abstract
The standard way to study Large Language Models (LLMs) through benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g. multiple choice questions). However, due to LLM’s highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts). We argue that context-dependence should be studied as another dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size. In this paper, we present a case-study about the stability of value expression over different contexts (simulated conversations on different topics), and as measured using a standard psychology questionnaire (PVQ) and a behavioral downstream task. We consider 19 open-sourced LLMs from five families. Reusing methods from psychology, we study Rank-order stability on the population (interpersonal) level, and Ipsative stability on the individual (intrapersonal) level. We explore two settings: with and without instructing LLMs to simulate particular personalities. We observe similar trends in the stability of models and model families - Mixtral, Mistral and Qwen families being more stable than LLaMa-2 and Phi - over those two settings, two different simulated populations, and even in the downstream behavioral task. When instructed to simulate particular personas, LLMs exhibit low Rank-Order stability, and this stability further diminishes with conversation length. This highlights the need for future research directions on LLMs that can coherently simulate a diversity of personas, as well as how context-dependence can be studied in more thorough and efficient ways. This paper provides a foundational step in that direction, and, to our knowledge, it is the first study of value stability in LLMs. The project website with code is available at https://sites.google.com/view/llmvaluestability.
Keywords: Large Language Models; LLM Stability; Context-dependence; Personal Values; Simulated populations
1 Introduction
In recent years, there has been an emergence of research using psychological tools to study Large Language Models (LLMs). In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas (?, ?). LLMs have also been used without providing such instructions, i.e. treated as a participant in a human study (?, ?). Questions in such studies have revolved around how Language Models express values (?, ?), personality traits (?, ?, ?), cognitive abilities (?, ?), and how they could replicate human data (?, ?).
Figure 1: How do LLM’s expressed values change as a function of context? An LLM is first prompted to play a specific role (e.g. Gandalf). Then, a conversation on a topic (e.g. joke) with an interlocutor model (same LLM prompted to simulate a human user) is generated. Then, the LLM simulating the persona is given a psychology questionnaire aimed to assess its expressed values. We study the stability of these expressed values across diverse conversation topics and lengths. We consider various personas to be simulated, as well as the case when the LLM is not prompted to play any particular persona. The messages and instructions in gray are set manually, and the messages in white are generated.
The use of psychological tools with LLMs opens up many scientific questions, for example regarding the nature of how the text generated by LLMs depends on context, i.e. information present in prompts or prior interaction with the model. Prompts or prior interaction, which we denote here as ’context’, can include any textual information such as instructions (e.g. personas to simulate), the dialogue history, stories written in specific styles, etc. Such contexts guide the generation of text by LLMs: different contexts may result in the expression of different behavior and values. This is sometimes beneficial and expected, e.g. an instruction to simulate some persona should influence the behavior and expressed values to be more aligned with that persona. However, this can also be problematic, e.g. a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways as we will highlight in experiments below. It should be noted that, depending on the application, different types of context-dependence can be beneficial or not.
The issue of unexpected context dependence in LLM is of crucial importance. Standard evaluation benchmarks, including those using psychological questionnaires to assess properties of LLMs, consist of sets of queries, often presented with a similar minimal context (e.g. knowledge or value-related questions presented as multiple choice questions with limited context). When deployed, LLMs are exposed to many new unforeseen contexts. This means that the standard benchmarks, by themselves, cannot estimate a model’s behavior in deployment (due to the LLMs’ highly context dependent nature). It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes in behavior.
This challenge is particularly acute with the use of psychological questionnaires aimed at measuring psychological dimensions like values. Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be significantly influenced by the content of a randomly picked Wikipedia article shown to them beforehand. As we will show below, such an assumption does not hold for many LLMs, and thus strongly limits general interpretability of using these questionnaires in a context-independent manner. It is thus key to understand better how LLMs’ behavior (e.g. expression of values) may maintain coherence or change as a function of various kinds of contexts (ranging from explicit instruction to play a particular persona to discussions about topics that seem unrelated to the expressed psychological dimensions one studies).
Previous research included certain experiments regarding unwanted context-based change (usually regarding syntactic changes in the prompt). These experiments led to conflicting results, sometimes showing robustness (?, ?, ?, ?, ?), and sometimes sensitivity (?, ?, ?). These inconclusive results motivate research into the nature and the extent of context-dependence of various LLMs.
In this paper, we present a case-study focusing on studying the stability of value expressed in 19 open-source LLMs from 5 families. First, we study to what extent can LLMs simulate various personas in coherent ways, i.e. expressed values should change according to the instructed persona, but not based on the topic of a conversation not related to values. In other words, we study to what extent do LLMs’ value profiles change in wanted ways (i.e. based on an instruction), while remaining robust to unwanted context-based change (i.e. based on different conversation topics). We instruct LLMs to simulate two populations: fictional characters and well-known real-world personas. In addition, we also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases. To our knowledge, this is the first study on value stability in LLMs.
We use the Schwartz’s theory of Basic Personal Values (?, ?) and the corresponding Portrait Values Questionnaire (PVQ-40) (?, ?). This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction). Following that research, we outline two types of value stability studies in the psychology literature: Rank-Order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level), see details below and on figures 2 and 3.
The main contributions of this paper are:
  • Introduction and adaptation of the methodology for evaluating Rank-Order and Ipsative stability in LLMs
  • First analysis of value stability across contexts (including various conversation lengths)
  • Systematic comparison of Rank-Order and Ipsative stability of 19 LLMs with and without instructing the models to simulate specific personas
  • Analysis of stability in a downstream behavioral task
Figure 2: Rank-Order stability An example of estimating Rank-Order stability for benevolence. In each context, characters are ordered according to their benevolence scores in that context. In this example, the orders are almost the same in contexts 1 and 2 (high Rank-Order stability), and very different in contexts 2 and 3 (low Rank-Order stability).
Figure 3: Ipsative stability An example of estimating Ipsative stability for a character (Gandalf). Values are ordered according to the character’s scores in each context. In this example, the orders are the same in contexts 1 and 2 (high Ipsative stability), and different in contexts 2 and 3 (low Ipsative stability).
2 Methods
This section discusses how we administer the PVQ questionnaire over different contexts to evaluate value stability. We conduct experiments in two ways: with and without instructing the models to simulate specific personas. Different contexts are induced by simulating conversations on different topics with a separate instance of the same model (the interlocutor). To set a conversation topic (e.g. joke) we manually set the first interlocutor’s message (e.g. "Tell me a joke."). We let the models exchange $n$ messages, manually set the last interlocutor’s message as the query (e.g. PVQ item), and record the model’s response. After the questionnaire was given in each context, we estimate the Ipsative and Rank-Order stability. This process is shown in figure 1.
2.1 Administering the questionnaire
Administering the questionnaire consists of the following steps depicted in figure 1:
1. A model is instructed to simulate a persona (optional). We study personas from two populations: 1) 60 fictional characters from J.R.R. Tolkien’s universe , and 2) 50 real world personas (see appendix A for details). A persona setting instruction (e.g. “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.”) is given to the model (see appendix A for details).
2. A separate interlocutor model instance is created. The interlocutor model is an instance of the same model as the one being tested. The interlocutor is given the following instruction: “You are simulating a human using a chatbot. Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be character_name.”
3. A conversation topic is induced. The first interlocutor’s message is manually set to induce one of the following topics: grammar, joke, poem, history, chess. For example, to induce the topic of “joke” it is set to “Tell me a joke.”. See appendix A for details.
4. A conversation is simulated. The two models are let to exchange $n$ messages. In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
5. A questionnaire is given. A questionnaire item is manually set as the interlocutor’s last message with a random order of suggested answers, and the model’s response is recorded. This is repeated for each question in parallel (with the same simulated conversation). That way, the model’s response is not influenced by responses to other questions.
6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See appendix A for details.
7. Stability is estimated. If a persona was provided in step 1, steps 1 to 6 are repeated for every persona in the simulated population. Then, the whole process is repeated with five random seeds. Stability is estimated for each seed and then averaged, i.e. value stability for one model is estimated from 50/60k queries, depending on the population (5 seeds x 5 topics x 50/60 personas x 40 PVQ items). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
If no persona was provided in step 1, steps 2 to 6 are repeated 50 times with different seeds. As no persona was provided, this process repeats the same experiment with 50 different permutations in the order of suggested answers, and therefore no additional seeds are needed. Ipsative stability is computed for each of the 50 permutations and then averaged, i.e. value stability for one model is estimated from 10-12k queries (5 topics x 50/60 permutations x 40 PVQ items).
2.2 Estimating the stability
We estimate two types of value stability: Rank-Order and Ipsative. Rank-Order estimates the stability of some value at the population (interpersonal) level, as the stability of the order of participants in expressing that value. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?". Ipsative stability estimates the stability at the individual (intrapersonal) level as the stability of individuals value hierarchies. Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?".
Rank-Order stability
Rank-Order stability is used to estimate the stability of some value inside a population. In psychology, Rank-Order stability for some value can be computed as the correlation in the order of individuals’ scores at two points in time (Spearman correlation between the participants’ ranks). Here, instead of comparing the participant ranks at two points in time, we compare it in different contexts (see Fig. 2). We evaluate a model in five different contexts and compute the stability for each pair of contexts, and estimate the final stability as the average of those pairs.
Ipsative (within-person) stability
Ipsative stability is used to estimate the stability of an individual’s value profile. In psychology, Ipsative stability can be computed as the correlation between the ranks of values for the same individual at two points in time (Spearman correlation between the values’ scores). Here, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (see Fig 3).
We evaluate models in five different contexts and compute the stability between each pair of contexts. We estimate the final stability by averaging over those pairs.
3 Experiments
This section provides an analysis of Ipsative and Rank-Order stability in LLMs. LLMs will be evaluated in two ways: with and without instructing the models to simulate particular personas. We aim to address the following questions:
  • How do different models and model families compare in terms of expressed value stability?
  • How does the stability of values expressed by LLMs compare to stability observed in human development?
  • Can LLMs keep coherent personas over longer conversations?
  • To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
3.1 Models
The LLaMa-2 (?, ?) family contains models with 7, 13 and 70 billion parameters. It also includes “chat” versions, which were fine-tuned on instructions and with RLHF (?, ?). The Mistral (?, ?) family contains base and instruction fine-tuned models with 7 billion parameters. Zephyr (?, ?) also belongs in this family as a DPO (?, ?) tuned version of the base Mistral model. The Mixtral (?, ?) family contains base and “instruct” models with 46.7 billion parameters. The “instruct” version was trained by supervised fine-tuning and DPO (?, ?). We consider these two models and their 4-bit quantized versions. The Phi (?, ?) family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters. From the Qwen (?, ?) model family we consider base models with 7, 14, and 74 billion parameters.
((a)) Rank-order value stability for LLMs simulating fictional characters.
((b)) Rank-order value stability for LLMs simulating real-world personas.
((c)) Rank-order donation stability on a downstream behavioral task for LLMs simulating fictional characters.
Figure 4: Rank-order stability ($Mean\pm SI(\alpha=0.05)$) of personal values and donations exhibited by LLMs following conversations on different topics. Consistent trends are visible: Mixtral, Qwen, and Mistral model families are the most stable, compared to LLaMa-2 and Phi families. All models exhibit lower than human stability, despite the comparison being skewed in their favor.
Figure 5: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs without the persona setting instructions. Mistral-Instruct-v0.1 and Qwen-72B models show the highest stability. Mixtral, Mistral and Qwen families are more stable. Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor.
3.2 How do different models and model families compare in terms of expressed value stability?
We evaluate the Rank-Order stability of LLMs instructed to simulate various personas. Figure 4 compares models’ value stability of two simulated populations: fictional characters (Fig. 3(a)) and real-world personas (Fig. 3(b)). On fictional characters the most stable model is Mixtral-Instruct-v1 ($r=0.42$), which is followed by its 4-bit quantized version ($r=0.32$), Mistral-Instruct-v2 ($r=0.31$), and Qwen-72B ($r=0.26$). A similar trend is observed on real-world personas, however, Qwen-72B ($r=0.5$) matches the stability of Mixtral-Instruct-v1 ($r=0.51$). More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral and Qwen families show more stability than LLaMa-2 and Phi.
Figure 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona. While similar trends of models are observed to those in the Rank-Order experiments, the models are less polarized. While Mixtral-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.8$), and Qwen-72B ($r=0.78$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-Instruct-v2 ($r=0.45$). Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat ($r=0.47$) and to a lesser extent in Phi-2 ($r=0.29$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.22$). The most stable model families are again Mixtral, Mistral and Qwen.
Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig. 5 is more stable than its base version. This effect is not as conclusive for Rank-order stability. As fine-tuning adapts the model towards instruction following, dialogues (chat textual format), and answering questions, it is expected to increase stability. However, it also often includes “aligning” the model by making it less prone to exhibit unwanted behavior, which can have a detrimental effect on simulating some personas such as villains. We hypothesize that this is the reason why we observe a consistent effect only on Ipsative stability.
3.3 How does the stability of values expressed by LLMs compare to stability observed in human development?
To get a more intuitive impression of the observed stability levels, we extract data from two longitudinal studies on humans. ? (?) followed 20-year-olds for eight years and ? (?) followed 10-year-old for 2 years (these changes are denoted by horizontal lines in figures 4 and 5). It is important to note that this comparison is skewed in the LLMs favor. It is easier for LLMs to show stability in the following ways: 1) human value changes were caused by much more drastic circumstances (years of development compared to topic change in LLMs) 2) the human population was more unstable (10-year-old and 20-year-olds compared to well-established fictional characters or real-world personas). Therefore, an argument can only be made in one direction: if some models show lower stability than that observed in humans, those models can be said to exhibit subhuman value stability.
Figures 3(a) and 3(b) show that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28). The fact that LLMs show lower stability despite the comparison being skewed in their favor shows that LLMs exhibit sub-human value stability and are significantly more susceptible to unexpected context changes. These results motivate reserach on LLMs focused on simulating populations.
Figure 5 shows the Ipsative stability of models that were not instructed to simulate a persona. Both Mixtral-Instruct models ($r=0.84$ and $r=0.8$), Qwen-72B ($r=0.78$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28)). Crucially, as discussed above, this does not imply that those models show human-level value stability, rather, the only insight is that other models show very low Ipsative stability.
Figure 6: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of the Mixtral-Instruct model simulating fictional characters. Stability decreases with longer simulated conversations.
Figure 7: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs with (Mixtral-Instruct) and without persona setting instructions. All models retain the same stability level in longer conversations.
3.4 Can LLMs keep coherent value profiles over longer conversations?
In the previous experiment, models were let to exchange $n=3$ messages (not including the manually set first and last interlocutor’s messages). Here, we evaluate the effect of simulated conversation length ($n$) on value stability.
Figure 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-Instruct model instructed to simulate fictional characters. Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-Instruct (the most stable model from Fig. 3(a)), and only on one population. We can see that, even for this most stable model, stability diminishes as conversations get longer. This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
Figure 7 shows the effect of conversation length on Ipsative stability. We compare the most stable models from Fig. 5 without persona instructions, and Mixtral-Instruct with instructions to simulate fictional characters. Ipsative stability remains stable regardless of the simulated conversation length for all models. Mixtral-Instruct with persona instruction ("Mixtral-Instruct (fict. char.)"), while highly stable, is less stable than uninstructed Mixtral-Instruct ("Mixtral-Intruct"). This implies that Mixtral-Instruct is slightly better adapted for use without the persona instructions.
Mixtral-Instruct with persona instructions exhibits a combination of low Rank-Order stability (Fig. 6) and high Ipsative stability (Fig. 7). This implies that the model is able to maintain characters consistently, however those characters might not be adequately impersonated (e.g. simulating different characters with similar value profiles). This hypothesis is confirmed in Appendix B.2. These results suggest that current LLMs are not well suited for use with persona setting instructions, and motivate future research on LLMs focused on simulating specific personas.
3.5 To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task (if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task). We construct a task where an LLMs can choose an amount of coins (0 to 10) to give a beggar. The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, hobbits). The average amount of donated coins is computed for each race. The stability of donated coins is then estimated in the same way as value stability, i.e. amounts donated to different races are treated in the same way as scores for different values. (see appendix A.6).
Figure 3(c) compares models according to their Rank-order stability on the donation task. The base Mixtral model (Mixtral-Base) shows the highest stability ($r=0.32$), and is closely followed by Mixtral-Instruct ($r=0.26$). Those models are then followed by Qwen-72B ($r=0.23$), the 4-bit quantized Mixtral base ($r=0.22$) and instruct ($r=0.18$) models, and then by Mistral-Instruct-v0.2 ($r=0.20$).
Compared to the results on PVQ (Fig. 3(a)), the overall trend of models is consistent with some differences. Qwen-72B, Mixtral-Instruct, and Mistral-Instruct-v0.2 are again among the most stable models. However, the Mixtral base models not only show stability, but even outperform their Mixtral-Instruct counterparts. Furthermore, zephyr-7b-beta shows lower, and the LLaMa-2-70b base model shows higher stability. The trends of model families are consistent: Mixtral, Mistral and Qwen are again the most stable, while Phi and LLaMa-2 show low stability. This experiment shows that, while the overall general trends are consistent between PVQ and the donation task, there are some differences (most notably the high stability of the base Mixtral model).
In appendix B.3 we study if behavior correlates with value expression, i.e. if donation amounts correlate positively with the expression of benevolence and universalism, and negatively with the expression of power and achievement.
4 Conclusion
This paper presents the first study into the stability of values expressed by Large Language Models. We consider (interpersonal) Rank-Order stability and (intrapersonal) Ipsative stability. We evaluate value stability over different contexts induced by simulating conversations about different topics. We conduct experiments with and without instructing the models to simulate particular personas. Over our experiments, we observed consistent trends of value stability: Mixtral, Mistral and Qwen model families were more stable (these trends are also confirmed on a downstream behavioral task). LLMs instructed to simulate personas exhibit much lower than human stability (despite the comparison being skewed in their favor), which further diminishes over longer conversations. This insight motivates future research on models specialized in simulating coherent populations of individuals.
This paper highlights how seemingly unrelated context changes can result in unpredictable and unwanted changes in behavior. We argue that context-dependence, and more precisely, value stability, should be seen as another dimension of LLM comparison alongside knowledge, model size, speed, and similar. Instead of evaluating LLMs with many different questions from a single minimal context, they should also be evaluated (in terms of their context-dependence and value stability) with the same questions asked in many different contexts. This study presents a first step in that direction.
Limitations Due to computational requirements for evaluating LLMs, the evaluation of stability considers only five different conversation topics and simulated rather short conversations (up to 9 messages). Greatly increasing the number of topics and conversation length could provide more precise insights into the stability of various models. For one model, Rank-Order value stability is estimated from 50/60k queries and Ipsative stability from 10/12k queries (depending on the simulated population). For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
Future work We believe that this paper opens many research avenues regarding context-dependence and value stability of LLMs. Similar questions to those explored in this paper could be explored for personality traits, cultural values, cognitive abilities and knowledge. An interesting direction is to explore if the same model can exhibit high stability in both settings with and without the persona instruction, or if specialized models are required. This paper opens a new area of research in creating, evaluating and analyzing models specialized in simulating coherent and diverse populations. Such models are needed for many applications such as replicating human studies (?, ?), simulating social interactions (?, ?), training teachers (?, ?), and many more.
5 Acknowledgments
Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151011996] made by GENCI. We would also like to thank Cédric Colas for many helpful discussions.
Appendix A Additional details on the methods
A.1 Constructing the simulated populations
We construct two simulated populations: Fictional characters and Real-world personas. Instructing a model to simulate a fictional or real persona that is well-known is a simple way to precisely define a persona’s character (assuming the models were trained on enough data related data). A list of personas from both populations is shown in table 1.
The Fictional characters population consists of characters from J.R.R. Tolkien’s universe. They were initially selected based on the length of their Wikipedia page (?, ?). Then, some characters were manually replaced by including more female characters and villains to make the population more balanced. Here is an example of an instruction inducing a persona from this population: “You are Gandalf from J. R. R. Tolkien’s Middle-earth legendarium.” This population contains a total of 60 fictional characters, which are listed in table 1.
The Real-world personas population contains personas from an online list of influential people (?, ?). Here is an example of an instruction inducing a persona from this population: “You are Marilyn Monroe (1926 – 1962) American actress, singer, model.” This population contains a total of 50 personas, which are listed in table 1.
A.2 Formatting the prompt
The full prompting procedure consists of five parts depicted by different colors in figure 11. The model is instructed to simulate a persona (blue), a conversation topic is induced (green), and a conversation is simulated (orange). The conversation is simulated using a separate instance of the same model: The Interlocutor model. This instance is given the instruction to simulate a human user (see section 2 and Fig. 1). Then, a query is posed (red), followed by a query string. The query string skews the next token distribution towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of relevant capital letters (e.g. A to F).
The precise prompt format depends on the LLM. We consider three types: 1) Base models: LLaMa-[7|13|70]b, Mi[s|x]tral-Base, phi-[1|2], Qwen-[7|14|72]B; 2) Tuned models with a system message input: LLaMa-[7|13|70]b-chat, zephyr-7b-beta; 2) Tuned models without a system message input: Mistral-Instruct-v0.[1|2], Mi[x]stral-Instruct. This specific system input is intended to provide more salient instructions to the model.
The base model format is depicted in figure 10(a). The persona is inducted to the Tested model with the following phrase: “CONTEXT: The following is a conversation with <persona>” (blue). Furthermore, the messages from the tested model are prefixed by “<persona_name>:” and from the interlocutor are prefixed by "USER:" (orange and purple). Hence, the tested model is modeling a conversation between a user and the persona. The Interlocutor model is prompted using the same format. It is given the following instruction: “CONTEXT: The following is a conversation between a human and a chatbot. The chatbot is pretending to be <persona_name>. The human’s every reply must be in one sentence only.”. The interlocutor model’s messages are prefixed by "HUMAN:", and the tested model’s responses by "<persona_name> (CHATBOT)". Hence, the interlocutor model is modeling a conversation between a human user and an LLM simulating a persona.
The formats for model tuned with and without the system message input are depicted in figure 10(b) and 10(c) respectively. Each tuned model has a specific template defining which parts of the text are messages from the user and which parts were generated by the model. For example, the Mistral model defines the user instruction between “[INST]” and “[/INST]” tags. In figures 10(b) and 10(c), keywords <system>, <user> and <assistant> denote the model specific tags. The only difference in format for the two types of models is the persona induction (blue). For models that have a system input, the persona inducing instruction (i.e. “You are <persona>”) is given through the system input. For models without such an input, the instruction is given as a user message followed by a manually set acknowledgment. 10(c) in blue). For both tuned model types, the same prompt format is used for the interlocutor model with the following instruction: "You are simulating a human using a chatbot. The chatbot is pretending to be <persona_name>. Your every reply must be in one sentence only."
A.3 Inducing a conversation topic
We induce a conversation topic by manually setting the first message of the interlocutor model. We consider the following five topics and their corresponding messages: 1) poem: “Hello, let’s write a poem together. You start by the first verse I’ll add the second one, and so on.”, 2) joke: “Tell me a joke.”, 3) history: “What is the significance of the battle of Hastings. Answer in two sentences.", 4) chess: "1. e4”, 5) grammar: “Can you check this sentence for grammar? Whilst Jane was waiting to meet hers friend their nose started bleeding.”.
A.4 Querying the model and evaluating its response
Following the simulated conversation, the model is presented with a query from a questionnaire or from the downstream donation task. The PVQ questionnaire is available in versions for male and female participants. We give the PVQ version according to the simulated persona’s gender. In experiments where no persona is simulated, and the questionnaire is given 50 times with different permutations in the order of suggested answer, male and female versions are used for 25 permutations each.
Following the full prompting procedure explained above and depicted in figure 11, a model generates a distribution for the next token. We score the probability of tokens corresponding to suggested answers. For the 6 point Likert scale, this corresponds to capital letters from A to F. The most probable letter is taken as the model’s answer.
A.5 Scoring the PVQ questionnaire
Once the answers have been collected, the scoring proceed as described in the questionnaires manual. The answers are coded from 1 to 5, where “Not like me at all” corresponds to 1 and “Very much like me” to 5. The per-participant mean score is subtracted from those scores to normalize the answers around 0. This step is crucial, as personal values are defined in relative terms (to what extent an individual values some value more than other values. The scores corresponding to each value are then averaged to obtain the per value scores.
A.6 Additional details on The Donation Task
The donation task was created using a template into which different fictional characters from the J.R.R. Tolkien’s universe were inserted. This task was inspired by a charity donation task in ? (?). Figure 12 shows an example of a query for an elf called Círdan, following a topic of “grammar”. A total of 100 different queries are constructed with five races (elves, dwarves, orcs, humans, and hobbits) each with 20 names (10 male and 10 female). Standard names for each race were chosen by taking random names from a list of all characters organized by race. We aimed to avoid specific characters for the beggars (e.g. we imply some unspecified hobbit called Jessamine). However, this could not be entirely achieved due to many names being unique.
We compute the average donated amount and estimate Rank-Order stability separately for each race separately, and then compute the final Rank-Order stability as the average over races. We treat donations to different races similarly to different values in PVQ. This choice was made because one could expect evil characters to donate more to orcs compared to other races, and vice versa. Hence, averaging over all characters might hide relevant trends in donation.
Appendix B Additional experiments and analyses
B.1 Visualization of Mixtral-Instruct value expression
In this section, we visualize the PVQ values expressed by the Mixtral-Instruct model along different contexts and seeds. We use PCA (?, ?) to visualize 250 dimensions (5 seeds x 5 contexts x 10 values) dimensions as two PCA components with $R^{2}=0.29$ and $R^{2}=0.12$ explained variance ratios. We used GPT-4 to classify characters into positive, neutral (more complex), and negative using the following prompt:
     Classify the following characters into positive/neutral/negative.    Create a two column table with the first column being the name of  the character and the second being the classification.    Gandalf  ...  ...  Gimli
Fig. 8 shows the representation of 60 fictional characters classified into positive, neutral (more complex) and negative characters. We can see that positive characters are grouped on the left, negative characters on the right, and neutral in the middle. This shows the Mixtral-Insturct model expressed values in a semantically plausible way.
Figure 8: PCA representation of different fictional characters simulated by the Mixtral-Instruct model. Positive characters (green) are grouped on the left side, negative characters (red) on the right side, and neutral (more complex) characters are in the middle.
B.2 Do simulated personas’ value profiles approach a neutral value profile with longer conversations?
In section 3.4 in the main text, we studied Mixtral-Instruct that was instructed to simulate fictional characters. We observed that Rank-order stability diminished and Ipsative stability stayed the same with longer conversations. This implied that simulated personas’ value profiles moved away from the instructed personas towards some neutral value profile. Here, we experimentally confirm this hypothesis by estimating the distance of simulated personas’ value profiles to an estimated neutral value profile. The distance is computed as Ipsative stability: correlation between the order of values in a simulated individual to those of the neutral value profile.
The neutral profile is estimated as follows. We evaluate the Mixtral-Instruct model without the persona setting instructions and without simulating a conversation (i.e. the questionnaire queries are given straight away). We repeat this process with 50 permutations in the order of suggested answers. To estimate the neutral profile, we average the value ranks over those permutations as shown in the following pseudocode:
     # value_score.shape == (50, 10)  value_scores = evaluate_model("Mixtral-Instruct", n_perm=50)  all_ranks = []  for i in range(50):      # permutation_value_profile.shape == (10)      permutation_scores = value_scores[i]      permuataion_value_ranks = compute_ranks(permuatation_scores)      all_ranks.append(permutation_values_ranks)    # all_ranks.shape == (50, 10)  # neutral_value_profile.shape == (10)  neutral_value_profile = all_rank.mean(axis=0)
Fig. 9 shows the similarity of simulated personas’ value profiles to the default profile, with the Rank-Order stability of simulated individuals between contexts (same as figure 3(a)). We can see that, as Rank-Order stability diminishes, simulated personas’ value profiles move closer to the default profile. This confirms our hypothesis that the diminishing Rank-Order stability is due to the model gradually “ignoring” the persona inducing instruction and moving all simulated personas’ value profiles closer to the neutral one.
Figure 9: Similarity of Mixtral-Instruct simulated fictional characters’ value profiles with the neutral value profile (blue) compared to the Rank-order stability. As conversations gets longer, simulated value profiles move away from the instructed persona toward a neutral one, resulting in lower Rank-order stability.
B.3 Is value expression correlated with behavior on a downstream task?
In section 3.5 in the main text, we studied if models that exhibit more stable value profiles also exhibit more stable behavior on a downstream task. Here, we study if value expression correlates with that behavior. We hypothesize that, for more stable models, simulated personas that exhibited higher universalism and benevolence will also donate more coins. Similarly, simulated personas that exhibited higher power and achievement should donate less.
We compute the correlations between the order of simulated participants in terms of expression of some value (e.g. Universalism) and the donation to each of the four fictional races (a total of 4 correlations), and compute the mean of those correlations. In doing so, the contexts are paired (e.g. the expression of Benevolence following a conversation about grammar is correlated with the amount donated to elves following a conversation about grammar).
Fig. 14 shows the correlation between rank-order of value expression on PVQ and the donation amount on a downstream task. As hypothesized, we can see that for most stable models: Mistral-Instruct-v1, Mixtral-Instruct (both versions), and Qwen-72B donations are correlated with Universalism and Benevolence, and negatively correlated with Power and Achievement. We again observe a trend in model families, with Mixtral, Mistral, and Qwen being more stable than LLaMa-2 and Phi. Having said that, neither model exhibited high correlation ($<0.3$). This experiment implies that, while expected positive and negative correlations between value expression and donation are observed, there is much room for improvement.
B.4 Do LLaMa-2 models exhibit low stability because of the persona induction method?
In the main text, LLaMa-2 chat models exhibited very low stability, but those models are also the only ones (apart from zephyr-7b-beta) which used the prompt template with the system message input as discussed in appendix A and shown in figure 10(b). Furthermore, Mistral-Instruct and Mixtral-Instruct, which showed high stability, used the template without the system message input (Fig. 10(c)). Therefore, we found it relevant to check that LLaMa-2 low stability is not caused by the prompting template but by the model itself.
Figure 13 shows the three LLaMa-2 chat tuned models with the two prompting templates. It compares inducing the persona through the system message (denoted by "*_sys"), as was done in the main text, to inducing it through the user message (denoted by "*_no_sys"), as was done for other models. We can see that neither prompt template enables the LLaMa-2 chat models to exhibit higher stability. This implies that the exhibited low stability is due to the models themselves, and not merely due to the choice of a prompting template.
Figure 10: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of chat-tuned LLaMa-2 models when the persona is induces through the system message (as was done in the main text) compared to the user message input (as was done for other models). LLaMa-2 models do not exhibit value stability in either setting. This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
((a)) Base models.
((b)) Instruction or Chat tuned model with the system input.
((c)) Instruction or Chat tuned model without the system input.
Figure 11: Prompt examples of administering a questionnaire to different models. For tuned models, <system>, <user>, and <assistant> are replaced with specific keywords defined by their fine-tuning. A persona (blue) and a conversation topic (green) are induced. A conversation is simulated (orange). A query from a questionnaire or a downstream task (purple) is given followed by the “query string’ (purple). The query string makes the next token distribution much more skewed towards capital letters denoting an answer. The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of capital letters from A to F.
Figure 12: An example of querying the model in the downstream donation task following a conversation about grammar. The example is the same as that in figure 10(b) with a different query (shown in red).
Table 1: A list of personas in the simulated populations
Population
Personas
Fictional characters
Gandalf, Gollum, Sméagol, Aragorn, Sauron, Saruman, Celeborn, Galadriel, Tom Bombadil, Elrond, Frodo Baggins, Maedhros, Finrod Felagund, Glorfindel, Goldberry, Bilbo Baggins, Smaug, Morgoth, Faramir, Éowyn, Samwise Gamgee, Fëanor, Théoden, Boromir, Túrin Turambar, Thranduil, Beorn, Arwen, Halbarad, Isildur, Gothmog (Balrog), Lungorthin, Celebrimbor, Gil-galad, Meriadoc Brandybuck, Treebeard, Shelob, Radagast, Elendil, Denethor, Éomer, Legolas, Húrin, Thorin Oakenshield, Peregrin Took, Thingol, Gríma Wormtongue, Eärendil, Elwing, Lúthien, Beren, Tuor, Idril, Finwë, Míriel, Ungoliant, Thuringwethil, Melian, Durin’s Bane, Gimli
Real-world persona
Marilyn Monroe, Abraham Lincoln, Nelson Mandela, Queen Elizabeth II, John F. Kennedy, Martin Luther King, Winston Churchill, Donald Trump, Bill Gates, Muhammad Ali, Mahatma Gandhi, Mother Teresa, Christopher Columbus, Charles Darwin, Elvis Presley, Albert Einstein, Paul McCartney, Queen Victoria, Pope Francis, Jawaharlal Nehru, Leonardo da Vinci, Vincent Van Gogh, Franklin D. Roosevelt, Pope John Paul II, Thomas Edison, Rosa Parks, Lyndon Johnson, Ludwig Beethoven, Oprah Winfrey, Indira Gandhi, Eva Peron, Benazir Bhutto, George Orwell, Vladimir Putin, Dalai Lama, Walt Disney, Neil Armstrong, Peter Sellers, Barack Obama, Malcolm X, J.K.Rowling, Richard Branson, Pele, Angelina Jolie, Jesse Owens, John Lennon, Henry Ford, Haile Selassie, Joseph Stalin, Lord Baden Powell
Figure 13: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of chat-tuned LLaMa-2 models when the persona is induces through the system message (as was done in the main text) compared to the user message input (as was done for other models). LLaMa-2 models do not exhibit value stability in either setting. This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
((a)) Universalism
((b)) Benevolence
((c)) Power
((d)) Achievement
Figure 14: Rank-order stability between value expression (on the PVQ questionnaire) and the donation amount (on a downstream task). For more stable modes, donations are correlated with Universalism and Benevolence and negatively correlated with power and achievement.

### v4

@@
 Abstract
-The standard way to study Large Language Models (LLMs) through benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g.
+The standard way to study Large Language Models (LLMs) with benchmarks or psychology questionnaires is to provide many different queries from similar minimal contexts (e.g.
 multiple choice questions).
-However, due to LLM’s highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts).
-We argue that context-dependence should be studied as another dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size.
-In this paper, we present a case-study about the stability of value expression over different contexts (simulated conversations on different topics), and as measured using a standard psychology questionnaire (PVQ) and a behavioral downstream task.
-We consider 19 open-sourced LLMs from five families.
+However, due to LLMs’ highly context-dependent nature, conclusions from such minimal-context evaluations may be little informative about the model’s behavior in deployment (where it will be exposed to many new contexts).
+We argue that context-dependence (specifically, value stability) should be studied as a specific property of LLMs and used as another dimension of LLM comparison (alongside others such as cognitive abilities, knowledge, or model size).
+We present a case-study on the stability of value expression over different contexts (simulated conversations on different topics) as measured using a standard psychology questionnaire (PVQ) and on behavioral downstream tasks.
 Reusing methods from psychology, we study Rank-order stability on the population (interpersonal) level, and Ipsative stability on the individual (intrapersonal) level.
-We explore two settings: with and without instructing LLMs to simulate particular personalities.
-We observe similar trends in the stability of models and model families - Mixtral, Mistral and Qwen families being more stable than LLaMa-2 and Phi - over those two settings, two different simulated populations, and even in the downstream behavioral task.
-When instructed to simulate particular personas, LLMs exhibit low Rank-Order stability, and this stability further diminishes with conversation length.
-This highlights the need for future research directions on LLMs that can coherently simulate a diversity of personas, as well as how context-dependence can be studied in more thorough and efficient ways.
+We consider two settings (with and without instructing LLMs to simulate particular personas), two simulated populations, and three downstream tasks.
+We observe consistent trends in the stability of models and model families - Mixtral, Mistral, GPT-3.5 and Qwen families are more stable than LLaMa-2 and Phi.
+The consistency of these trends implies that some models exhibit higher value-stability than others, and that value stability can be estimated with the set of introduced methodological tools.
+When instructed to simulate particular personas, LLMs exhibit low Rank-order stability, which further diminishes with conversation length.
+This highlights the need for future research on LLMs that coherently simulate different personas.
 This paper provides a foundational step in that direction, and, to our knowledge, it is the first study of value stability in LLMs.
 The project website with code is available at https://sites.google.com/view/llmvaluestability.
-Keywords: Large Language Models; LLM Stability; Context-dependence; Personal Values; Simulated populations
+Keywords: Large Language Models; LLM Stability; Value-stability, Basic Personal Values, Role-play, Context-dependence; Simulated populations
 1 Introduction
 In recent years, there has been an emergence of research using psychological tools to study Large Language Models (LLMs).
-In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas (?, ?).
+In those studies, LLMs have often been used to simulate populations by instructing them to simulate different personas [1].
 LLMs have also been used without providing such instructions, i.e.
-treated as a participant in a human study (?, ?).
-Questions in such studies have revolved around how Language Models express values (?, ?), personality traits (?, ?, ?), cognitive abilities (?, ?), and how they could replicate human data (?, ?).
-Figure 1: How do LLM’s expressed values change as a function of context?
-An LLM is first prompted to play a specific role (e.g.
+treated as a participant in a human study [2].
+Questions in such studies have revolved around how Language Models express values [3], personality traits [4, 5], cognitive abilities [6], and how they could replicate human data [7].
+Figure 1: Evaluating the expressed value profile in context The tested LLM is prompted to play a specific role (e.g.
 Gandalf).
-Then, a conversation on a topic (e.g.
-joke) with an interlocutor model (same LLM prompted to simulate a human user) is generated.
-Then, the LLM simulating the persona is given a psychology questionnaire aimed to assess its expressed values.
-We study the stability of these expressed values across diverse conversation topics and lengths.
-We consider various personas to be simulated, as well as the case when the LLM is not prompted to play any particular persona.
+We simulate a conversation on a topic (e.g.
+joke) with an interlocutor model (same LLM prompted to simulate a human user).
+Then, the tested LLM is given a psychology questionnaire (PVQ-40) aimed to assess its expressed values.
+We study the stability of these expressed values (and of behavior on downstream tasks) across diverse conversation topics and lengths.
+We consider the simulation of various fictional and real-world personas, as well as the case when the LLM is not prompted to play any particular persona.
 The messages and instructions in gray are set manually, and the messages in white are generated.
-The use of psychological tools with LLMs opens up many scientific questions, for example regarding the nature of how the text generated by LLMs depends on context, i.e.
-information present in prompts or prior interaction with the model.
+The use of psychological tools with LLMs opens up many scientific questions, for instance regarding the nature of how the LLM-generated text depends on context, i.e.
+information present in the prompt or prior interaction with the model.
 Prompts or prior interaction, which we denote here as ’context’, can include any textual information such as instructions (e.g.
@@
 However, this can also be problematic, e.g.
-a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways as we will highlight in experiments below.
+a specific conversation topic may drastically influence the expressed behavior and values in unexpected ways, as we will highlight in experiments below.
 It should be noted that, depending on the application, different types of context-dependence can be beneficial or not.
+For example, to avoid malicious misuse of an AI-assistant, it might be beneficial to not express different values even when instructed to simulate different personas.
+On the other hand, simulating diverse populations of agents [8] or humans [7, 1] requires exactly this kind of context-dependence.
 The issue of unexpected context dependence in LLM is of crucial importance.
@@
 This means that the standard benchmarks, by themselves, cannot estimate a model’s behavior in deployment (due to the LLMs’ highly context dependent nature).
-It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes in behavior.
+It is therefore crucial to evaluate the robustness of different models to unexpected context-based changes as a property, which can then be used as a dimension of LLM comparison alongside others such as cognitive abilities, knowledge, or model size.
 This challenge is particularly acute with the use of psychological questionnaires aimed at measuring psychological dimensions like values.
-Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be significantly influenced by the content of a randomly picked Wikipedia article shown to them beforehand.
+Those tools were initially designed to probe humans, and thus make various assumptions about humans: for example, it is expected that the answers of most humans to questionnaires about value preferences should not be drastically influenced by the content of a random conversation.
 As we will show below, such an assumption does not hold for many LLMs, and thus strongly limits general interpretability of using these questionnaires in a context-independent manner.
@@
 Previous research included certain experiments regarding unwanted context-based change (usually regarding syntactic changes in the prompt).
-These experiments led to conflicting results, sometimes showing robustness (?, ?, ?, ?, ?), and sometimes sensitivity (?, ?, ?).
+These experiments led to conflicting results, sometimes showing robustness [9, 10, 4, 11], and sometimes sensitivity [2, 11].
 These inconclusive results motivate research into the nature and the extent of context-dependence of various LLMs.
-In this paper, we present a case-study focusing on studying the stability of value expressed in 19 open-source LLMs from 5 families.
-First, we study to what extent can LLMs simulate various personas in coherent ways, i.e.
+In this paper, we present a case-study focusing on studying the stability of value expressed in 21 LLMs from 6 families.
+We simulate conversations with another instance of the LLM, which was instructed to simulate a human user (see Fig.
+1).
+We study to what extent can LLMs simulate various personas in coherent ways, i.e.
 expressed values should change according to the instructed persona, but not based on the topic of a conversation not related to values.
-In other words, we study to what extent do LLMs’ value profiles change in wanted ways (i.e.
+In other words, we study to what extent do LLM-expressed value profiles change in wanted ways (i.e.
 based on an instruction), while remaining robust to unwanted context-based change (i.e.
 based on different conversation topics).
-We instruct LLMs to simulate two populations: fictional characters and well-known real-world personas.
-In addition, we also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases.
+We instruct LLMs to simulate various fictional characters and real-world personas (from different countries and cultural backgrounds).
+We also study the coherence and robustness of LLMs’ value expression when they are not instructed to simulate any specific persona, corresponding to frequent real-world use cases.
 To our knowledge, this is the first study on value stability in LLMs.
-We use the Schwartz’s theory of Basic Personal Values (?, ?) and the corresponding Portrait Values Questionnaire (PVQ-40) (?, ?).
-This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction).
-Following that research, we outline two types of value stability studies in the psychology literature: Rank-Order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level), see details below and on figures 2 and 3.
-The main contributions of this paper are:
-  • Introduction and adaptation of the methodology for evaluating Rank-Order and Ipsative stability in LLMs
-  • First analysis of value stability across contexts (including various conversation lengths)
-  • Systematic comparison of Rank-Order and Ipsative stability of 19 LLMs with and without instructing the models to simulate specific personas
-  • Analysis of stability in a downstream behavioral task
-Figure 2: Rank-Order stability An example of estimating Rank-Order stability for benevolence.
+We use the Schwartz’s theory of Basic Personal Values [12, 13] and the corresponding Portrait Values Questionnaire (PVQ-40) [14].
+This theory and questionnaire has been thoroughly studied and validated in the field of psychology, and it outlines universal ten basic personal values (Universalism, Benevolence, Conformity, Tradition, Security, Power, Achievement, Hedonism, Stimulation, and Self-Direction).
+The PVQ-40 questionnaire has been used to study values in a large diversity of countries and cultures [15].
+Refer to Appendix A for details.
+We outline two types of value stability commonly studied in the field of psychology: Rank-order (on a population/interpersonal level) and Ipsative (on an individual/intrapersonal level) (see Figures 2 and 3).
+Figure 2: Rank-order stability An example of estimating Rank-order stability of benevolence.
 In each context, characters are ordered according to their benevolence scores in that context.
 In this example, the orders are almost the same in contexts 1 and 2 (high Rank-order stability), and very different in contexts 2 and 3 (low Rank-order stability).
-Figure 3: Ipsative stability An example of estimating Ipsative stability for a character (Gandalf).
+Figure 3: Ipsative stability An example of estimating Ipsative stability for a fictional character (Gandalf).
 Values are ordered according to the character’s scores in each context.
 In this example, the orders are the same in contexts 1 and 2 (high Ipsative stability), and different in contexts 2 and 3 (low Ipsative stability).
+The main contributions of this paper are:
+  • Conceptualization of context-dependence (specifically, value-stability) as a property of LLM, to be used as another dimension of LLM comparison alongside others such as cognitive abilities, model size or inference speed.
+  • A set of methodological tools enabling the study of (intrapersonal and interpersonal) value stability and context-dependence in LLMs
+  • First analysis of stability of value expression in LLMs across contexts.
+The analysis was conducted over: 21 LLMs from six families, two types of stability, two settings (with and without instructing them to simulate specific personas), two simulated populations, and three downstream tasks.
+2 Related Work
+There has been a growing number of works using psychological questionnaires to study LLMs.
+[2] evaluated GPT-3 on a battery of vignette-based cognitive tests, and [6] evaluated a LaMDa model on a battery of classical developmental tasks and compared its performance to that of human children.
+Some works have evaluated LLMs using personality questionnaires [5, 16] and tests for creativity [17] Although not directly using psychological questionnaires, there are works heavily inspired by psychology which estimate LLMs’ Theory of Mind through textual representations of standard False-belief tasks [18, 19].
+Likewise, psychological questionnaires have been used with LLMs to inform psychological [20] and neuroscience [21, 22] hypotheses.
+A body of work has studied LMs as simulating a diversity of individuals and cultures [23].
+[24] train LSTM-based [25] conversational models to simulate different personas.
+[26] introduced the metaphor of role-playing, where an LLM always chooses a character to role-play based on context.
+Cultural expression has been studied by inducing personalities from different countries [27] and by querying the model in different languages [28].
+[29] also used those methods to compare LLM responses to human data from the World Values Survey [30] and the Pew Global Attitudes Survey [31].
+[10] induced personas of various demographic groups and observed a left-leaning bias.
+[32] fine-tune LLM to simulate different high- and low-resource cultures.
+[33] study the cultural evolution of stories in populations of LLMs induced with different personalities.
+[34] simulate children of different ages to recreate developmental stages of exploration, and simulate different domain experts to improve question answering.
+[35] show that inducing famous personalities crease toxicity.
+[7] replicate psychological studies with humans by varying the participants’ names to simulate a culturally diverse population.
+Similarly, [1] replicate data from human studies by prompting the model with backstories of real human participants in those original studies.
+LLMs have also been used to simulate students in order to train teachers [36].
+Previous work has also highlighted the influence of context on LLMs behavior [37].
+[38] show that exposing an LLM to some statement increases its perceived truthfulness at a later time, and [39] demonstrate the tendency of models to repeat back the user’s answer.
+[40] highlights the need for multi-turn evaluation to uncover attractors in the cultural evolution of stories generated by LLMs.
+The most similar work to ours is a concurrent paper studying the coherence of simulated personas in general (as opposed to our specific focus on personal value stability in simulated individuals and populations).
+That work proposes to increase the similarity of LLaMa-2-70b-chat [41] model’s answers before and after simulated conversations by reweighting the instruction’s attention weights [42].
 3 Methods
@@
 After the questionnaire was given in each context, we estimate the Ipsative and Rank-order stability.
-This process is shown in figure 1.
+This process is shown on Fig 1.
 3.1 Administering the questionnaire
-Administering the questionnaire consists of the following steps depicted in figure 1:
+Administering the questionnaire consists of the following steps depicted in Fig 1:
 1.
@@
 We study personas from two populations: 1) 60 fictional characters from J.R.R.
-Tolkien’s universe , and 2) 50 real world personas (see appendix A for details).
+Tolkien’s universe , and 2) 50 real-world personas (see Appendix B.1 for details).
+Instructing a model to simulate a well-known persona is a simple way to precisely define a persona’s character (assuming the models were trained on enough of related data).
 A persona setting instruction (e.g.
@@
 R.
-Tolkien’s Middle-earth legendarium.”) is given to the model (see appendix A for details).
+Tolkien’s Middle-earth legendarium.”) is given to the model (see Appendix B.2 for details).
 2.
@@
 The interlocutor is given the following instruction: “You are simulating a human using a chatbot.
-Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be character_name.”
+Your every reply must be in one sentence only.” If a persona was provided in step 1, the following sentence is added as the second sentence: “The chatbot is pretending to be <character_name>.”
 3.
@@
 For example, to induce the topic of “joke” it is set to “Tell me a joke.”.
-See appendix A for details.
-4. A conversation is simulated. The two models are let to exchange $n$ messages. In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
+See Appendix B.3 for details.
+4.
+A conversation is simulated.
+The two models are let to exchange $n$ messages.
+[^1] The default parameters from each model’s card on the huggingface website (https://huggingface.co/models) are used.
+In our experiments, $n$ is set to 3 (except when studying the effect of $n$ on stability).
 5.
 A questionnaire is given.
-A questionnaire item is manually set as the interlocutor’s last message with a random order of suggested answers, and the model’s response is recorded.
+A questionnaire item is manually set as the interlocutor’s last message (with a random order of suggested answers), and the model’s response is recorded.
 This is repeated for each question in parallel (with the same simulated conversation).
-That way, the model’s response is not influenced by responses to other questions.
-6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See appendix A for details.
+That way, the model’s response is not influenced by the responses to other questions.
+6. A questionnaire is scored. The responses are scored to obtain the scores for the ten values. See Appendix B.5 for details.
 7.
@@
 value stability for one model is estimated from 50/60k queries, depending on the population (5 seeds x 5 topics x 50/60 personas x 40 PVQ items).
-For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
+For reference, MMLU [43] (a commonly used general knowledge benchmark) contains 14k test questions.
 If no persona was provided in step 1, steps 2 to 6 are repeated 50 times with different seeds.
@@
 We estimate two types of value stability: Rank-order and Ipsative.
-Rank-Order estimates the stability of some value at the population (interpersonal) level, as the stability of the order of participants in expressing that value.
-Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?".
-Ipsative stability estimates the stability at the individual (intrapersonal) level as the stability of individuals value hierarchies.
-Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?".
+Rank-order estimates the stability of some value at the population (interpersonal) level, and Ipsative stability estimates the stability of value profiles at the individual (intrapersonal) level.
 Rank-order stability
 Rank-order stability is used to estimate the stability of some value inside a population.
-In psychology, Rank-Order stability for some value can be computed as the correlation in the order of individuals’ scores at two points in time (Spearman correlation between the participants’ ranks).
-Here, instead of comparing the participant ranks at two points in time, we compare it in different contexts (see Fig.
-2).
-We evaluate a model in five different contexts and compute the stability for each pair of contexts, and estimate the final stability as the average of those pairs.
+In psychology, it is computed as the correlation in the order of individuals at two points in time (individuals are ordered based on their expression of that value).
+Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Jane does?".
+As shown in Fig 2, instead of comparing two points in time, we compare different contexts (simulated conversations of different topics).
 Ipsative (within-person) stability
 Ipsative stability is used to estimate the stability of an individual’s value profile.
-In psychology, Ipsative stability can be computed as the correlation between the ranks of values for the same individual at two points in time (Spearman correlation between the values’ scores).
-Here, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (see Fig 3).
-We evaluate models in five different contexts and compute the stability between each pair of contexts. We estimate the final stability by averaging over those pairs.
+In psychology, it is computed as the correlation in the order of values expressed by the same individual at two points in time.
+Intuitively, this can be seen as addressing the following question: "Does Jack always value Tradition more than Benevolence?".
+As shown in Fig 3, instead of evaluating the value profile at two points in time, we evaluate it in different contexts (simulated conversations of different topics).
+For both Rank-order and Ipsative stability, we evaluate models in five different contexts, compute the stability between each pair of contexts, and estimate the final stability by averaging over those pairs.
+For Rank-order stability, we additionally average the stabilities of the ten values.
 4 Experiments
-This section provides an analysis of Ipsative and Rank-Order stability in LLMs.
-LLMs will be evaluated in two ways: with and without instructing the models to simulate particular personas.
+This section provides an analysis of context-dependence (specifically, value-stability) as a property of LLMs.
+We observe consistent trends along different experimental settings, implying that some models and model families exhibit more favorable context-dependence compared to others.
+LLMs are evaluated in two ways: with and without instructing them to simulate particular personas.
 We aim to address the following questions:
   • How do different models and model families compare in terms of expressed value stability?
-  • How does the stability of values expressed by LLMs compare to stability observed in human development?
+  • How does the stability of values expressed by LLMs compare to that observed in human development?
   • Can LLMs keep coherent personas over longer conversations?
-  • To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
+  • To what extent do conclusions made with PVQ transfer to downstream behavioral tasks?
+  • Is value expression on PVQ correlated with behavior on a downstream task?
+  • How additional contexts affect stability estimates?
 4.1 Models
-The LLaMa-2 (?, ?) family contains models with 7, 13 and 70 billion parameters.
-It also includes “chat” versions, which were fine-tuned on instructions and with RLHF (?, ?).
-The Mistral (?, ?) family contains base and instruction fine-tuned models with 7 billion parameters.
-Zephyr (?, ?) also belongs in this family as a DPO (?, ?) tuned version of the base Mistral model.
-The Mixtral (?, ?) family contains base and “instruct” models with 46.7 billion parameters.
-The “instruct” version was trained by supervised fine-tuning and DPO (?, ?).
+We consider 21 LLMs from the following six model families: LLaMa-2 [41], Mistral [44], Mixtral [45], Phi [46], Qwen [47], and GPT-3.5 [48].
+The LLaMa-2 family contains models with 7, 13 and 70 billion parameters (“llama_2_[7|13|70]b”) trained with 2T tokens.
+It also includes “chat” versions (“llama_2_[7|13|70]b_chat”), which were fine-tuned on instructions and with RLHF [49].
+The Mistral family contains base (“Mistral-7B-v0.1”) and instruction fine-tuned models (“Mistral-7B-Instruct-v0.[1|2]”) with 7 billion parameters.
+Zephyr [50] (“zephyr-7b-beta”) also belongs in this family as a DPO [51] tuned version of the base Mistral model.
+The Mixtral family contains base (“Mixtral-8x7B-v0.1”) and “instruct” (“Mixtral-8x7B-Instruct-v0.1”) models with 46.7 billion parameters.
+Those are Mixture-of-Experts models, which means that only 12.9 billion parameters are used per token.
+The “instruct” version was trained by supervised fine-tuning and DPO.
 We consider these two models and their 4-bit quantized versions.
-The Phi (?, ?) family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters.
-From the Qwen (?, ?) model family we consider base models with 7, 14, and 74 billion parameters.
-((a)) Rank-order value stability for LLMs simulating fictional characters.
-((b)) Rank-order value stability for LLMs simulating real-world personas.
-((c)) Rank-order donation stability on a downstream behavioral task for LLMs simulating fictional characters.
-Figure 4: Rank-order stability ($Mean\pm SI(\alpha=0.05)$) of personal values and donations exhibited by LLMs following conversations on different topics.
-Consistent trends are visible: Mixtral, Qwen, and Mistral model families are the most stable, compared to LLaMa-2 and Phi families.
+The Phi family contains smaller base models, of which we consider two models with 1.3 and 2.7 billion parameters (“phi-[1|2]”).
+From the Qwen family we consider base models with 7, 14, and 74 billion parameters (“Qwen-[7|14 |74]B”), which were trained on 2.2T 2.4T 3T tokens, respectively.
+From the GPT-3.5 family, we consider the latest two versions: from January 2024 (“gpt-3.5-turbo-0125”) and from October 2023 (“gpt-3.5-turbo-1106”).
+4.2 How do different models and model families compare in terms of expressed value stability?
+Rank-order stability
+We evaluate the Rank-order stability of LLMs instructed to simulate various personas.
+Fig 4 compares models’ value stability of two simulated populations: fictional characters (Fig 4(a)) and real-world personas (Fig 4(b)).
+Statistical analyses for Figs 4(a) and 4(b) are shown in Appendix E in Figs 16 and 17, respectively.
+On fictional characters (Fig 4(a)) the most stable model is Mixtral-8x7B-Instruct-v0.1 ($r=0.43$), which is followed by its 4-bit quantized version ($r=0.3$), Mistral-7B-Instruct-v0.2 ($r=0.28$), Qwen-72B ($r=0.24$), gpt-3.5-turbo-1106 ($r=0.20$), and gpt-3.5-turbo-0125 ($r=0.15$).
+A similar trend is observed on real-world personas (Fig 4(a)), however, Qwen-72B ($r=0.46$) approaches the stability of Mixtral-8x7B-Instruct-v0.1 ($r=0.5$).
+More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral, GPT-3.5, and Qwen families show more stability than LLaMa-2 and Phi.
+((a)) Fictional characters simulated population
+((b)) Real-world personas simulated population
+Figure 4: Rank-order stability with PVQ Rank-order stability ($Mean\pm SE$) of personal values (PVQ) exhibited by simulated participants (fictional characters or real-world personas) following conversations on different topics (correlation of simulated participants’ value expression in different contexts).
+Consistent trends are visible: Mixtral, Qwen, Mistral, and GPT-3.5 model families are more stable than LLaMa-2 and Phi families.
 All models exhibit lower than human stability, despite the comparison being skewed in their favor.
-Figure 5: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs without the persona setting instructions.
-Mistral-Instruct-v0.1 and Qwen-72B models show the highest stability.
-Mixtral, Mistral and Qwen families are more stable.
-Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor.
-3.2 How do different models and model families compare in terms of expressed value stability?
-We evaluate the Rank-Order stability of LLMs instructed to simulate various personas.
-Figure 4 compares models’ value stability of two simulated populations: fictional characters (Fig.
-3(a)) and real-world personas (Fig.
-3(b)).
-On fictional characters the most stable model is Mixtral-Instruct-v1 ($r=0.42$), which is followed by its 4-bit quantized version ($r=0.32$), Mistral-Instruct-v2 ($r=0.31$), and Qwen-72B ($r=0.26$).
-A similar trend is observed on real-world personas, however, Qwen-72B ($r=0.5$) matches the stability of Mixtral-Instruct-v1 ($r=0.51$).
-More generally, we observe consistent trends in terms of model families in both simulated populations: Mixtral, Mistral and Qwen families show more stability than LLaMa-2 and Phi.
-Figure 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona.
+LLMs are simulating two populations: (a) fictional characters, and (b) real-world personas.
+For statistical tests, refer to Appendix E (Figs 17 and 16)
+Ipsative stability
+Fig 5 compares the Ipsative stability of LLMs without instructing them to simulate any particular persona.
+The statistical analysis is shown in Appendix E in Fig 18.
 While similar trends of models are observed to those in the Rank-order experiments, the models are less polarized.
-While Mixtral-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.8$), and Qwen-72B ($r=0.78$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-Instruct-v2 ($r=0.45$).
-Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat ($r=0.47$) and to a lesser extent in Phi-2 ($r=0.29$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.22$).
-The most stable model families are again Mixtral, Mistral and Qwen.
-Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig.
-5 is more stable than its base version.
+While Mixtral-8x7B-Instruct-v0.1 ($r=0.84$), its 4-bit quantized version ($r=0.82$), and Qwen-72B ($r=0.73$) are again the most stable models, zephyr-7b-beta ($r=0.62$) is more stable than Mistral-7B-Instruct-v0.2 ($r=0.48$).
+Furthermore, compared to the previous experiment, stability is also observed in LLaMa-2-70b-chat($r=0.47$) and to a lesser extent in Phi-2 ($r=0.3$), LLaMa-2-70b ($r=0.17$), and Qwen-7B ($r=0.18$).
+The most stable model families are again Mixtral, Mistral, GPT-3.5 and Qwen.
+Instruction or chat fine-tuning seems to be beneficial for Ipsative stability, as every tuned model in Fig 5 is more stable than its base version.
 This effect is not as conclusive for Rank-order stability.
@@
 We hypothesize that this is the reason why we observe a consistent effect only on Ipsative stability.
+Figure 5: Ipsative stability with PVQ Ipsative stability ($Mean\pm SE$) of personal values (PVQ) exhibited by LLMs without the persona setting instructions (correlation of value hierarchies in different contexts).
+Mistral-7B-Instruct-v0.1 and Qwen-72B models show the highest stability.
+Mixtral, Mistral, Qwen and GPT-3.5 families are more stable.
+Human change is shown for reference, but no strong conclusions can be made because the comparison is skewed in the LLMs’ favor.
+For statistical tests, refer to Appendix E (Fig 18).
 4.3 How does the stability of values expressed by LLMs compare to stability observed in human development?
 To get a more intuitive impression of the observed stability levels, we extract data from two longitudinal studies on humans.
-?
-(?) followed 20-year-olds for eight years and ?
-(?) followed 10-year-old for 2 years (these changes are denoted by horizontal lines in figures 4 and 5).
+[52] followed 20-year-olds for eight years and [53] followed 10-year-olds for 2 years (these changes are denoted by horizontal lines in Fig 4 and Fig 5).
 It is important to note that this comparison is skewed in the LLMs favor.
@@
 Therefore, an argument can only be made in one direction: if some models show lower stability than that observed in humans, those models can be said to exhibit subhuman value stability.
-Figures 3(a) and 3(b) show that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28).
+Fig 4 shows that all models, when instructed to simulate various personas, exhibit much lower Rank-order stability than that observed in human populations ($r=0.57$ for ages 10 to 12, and $r=0.66$ for ages 20 to 28).
 The fact that LLMs show lower stability despite the comparison being skewed in their favor shows that LLMs exhibit sub-human value stability and are significantly more susceptible to unexpected context changes.
-These results motivate reserach on LLMs focused on simulating populations.
-Figure 5 shows the Ipsative stability of models that were not instructed to simulate a persona.
-Both Mixtral-Instruct models ($r=0.84$ and $r=0.8$), Qwen-72B ($r=0.78$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28)).
+These results motivate research on LLMs focused on simulating populations.
+Fig 5 shows the Ipsative stability of models that were not instructed to simulate a persona.
+Both Mixtral-8x7B-Instruct-v0.1 models ($r=0.84$ and $r=0.82$), Qwen-72B ($r=0.73$), and zephyr-7b-beta ($r=0.62$) do not exhibit lower stability than that observed in humans ($r=0.66$ for ages 10 to 12, and $r=0.59$ for ages 20 to 28).
 Crucially, as discussed above, this does not imply that those models show human-level value stability, rather, the only insight is that other models show very low Ipsative stability.
-Figure 6: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of the Mixtral-Instruct model simulating fictional characters. Stability decreases with longer simulated conversations.
-Figure 7: Ipsative value stability ($Mean\pm SI(\alpha=0.05)$) of LLMs with (Mixtral-Instruct) and without persona setting instructions.
+4.4 Can LLMs keep coherent value profiles over longer conversations?
+In the previous experiment, models were let to exchange $n=3$ messages (not counting the manually set first and last interlocutor’s messages).
+Here, we evaluate the effect of increased simulated conversation length ($n$) on value stability.
+Fig 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-8x7B-Instruct-v0.1 model instructed to simulate fictional characters.
+Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-8x7B-Instruct-v0.1 (the most stable model from Fig 4), and only on one population (fictional characters).
+We can see that, even for this most stable model, stability diminishes as conversations get longer.
+It gradually diminishes from $r=0.42$ $(n=3)$ to $r=0.15$ $(n=43)$.
+We can also see that the stability seems to converge after 35 messages with only a slight drop from 35 to 43 simulated messages ($r=0.166$ to $r=0.162$).
+This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
+Figure 6: Rank-order stability with longer conversations Rank-order value stability ($Mean\pm SE$) following conversations of different length for the Mixtral-8x7B-Instruct-v0.1 model simulating fictional characters (correlation of simulated participants’ value expression in different contexts).
+Stability decreases with longer simulated conversations.
+Ipsative stability
+Fig 7 shows the effect of conversation length on Ipsative stability.
+We compare the most stable models from Fig 5 without persona instructions, and Mixtral-8x7B-Instruct-v0.1 with instructions to simulate fictional characters.
+Ipsative stability remains stable regardless of the simulated conversation length for all models.
+Mixtral-8x7B-Instruct-v0.1 with persona instructions (“Mixtral-8x7B-Instruct-v0.1 (fict.
+char.)”), while highly stable, is less stable than the uninstructed model (“Mixtral-8x7B-Instruct-v0.1”).
+This implies that Mixtral-8x7B-Instruct-v0.1 is slightly better adapted for use without the persona instructions.
+Figure 7: Ipsative stability with longer conversations Ipsative value stability ($Mean\pm SE$) of LLMs with (fict.
+char.) and without persona setting instructions (correlation of value hierarchies in different contexts).
 All models retain the same stability level in longer conversations.
-3.4 Can LLMs keep coherent value profiles over longer conversations?
-In the previous experiment, models were let to exchange $n=3$ messages (not including the manually set first and last interlocutor’s messages).
-Here, we evaluate the effect of simulated conversation length ($n$) on value stability.
-Figure 6 shows the effect of simulated conversation length on Rank-order stability expressed by the Mixtral-Instruct model instructed to simulate fictional characters.
-Due to computational constraints (evaluating a model on one population requires 300k queries), we conduct this experiment only with Mixtral-Instruct (the most stable model from Fig.
-3(a)), and only on one population.
-We can see that, even for this most stable model, stability diminishes as conversations get longer.
-This experiment highlights the limitations of LLMs in maintaining coherent interpersonal value profiles over longer conversations.
-Figure 7 shows the effect of conversation length on Ipsative stability.
-We compare the most stable models from Fig.
-5 without persona instructions, and Mixtral-Instruct with instructions to simulate fictional characters.
-Ipsative stability remains stable regardless of the simulated conversation length for all models.
-Mixtral-Instruct with persona instruction ("Mixtral-Instruct (fict.
-char.)"), while highly stable, is less stable than uninstructed Mixtral-Instruct ("Mixtral-Intruct").
-This implies that Mixtral-Instruct is slightly better adapted for use without the persona instructions.
-Mixtral-Instruct with persona instructions exhibits a combination of low Rank-Order stability (Fig.
-6) and high Ipsative stability (Fig.
-7).
-This implies that the model is able to maintain characters consistently, however those characters might not be adequately impersonated (e.g.
-simulating different characters with similar value profiles).
-This hypothesis is confirmed in Appendix B.2.
+Mixtral-8x7B-Instruct-v0.1 with persona instructions exhibits a combination of decreasing low Rank-order stability (Fig 6) and high Ipsative stability (Fig 7).
+This implies that as conversations get longer, the simulated persona drifts towards a similar value profile.
+This hypothesis is confirmed in Appendix D.2.
 These results suggest that current LLMs are not well suited for use with persona setting instructions, and motivate future research on LLMs focused on simulating specific personas.
-3.5 To what extent do conclusions made with PVQ transfer to a downstream behavioral task?
-In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task (if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task).
-We construct a task where an LLMs can choose an amount of coins (0 to 10) to give a beggar.
-The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, hobbits).
+We hypothesize this to be a consequence of instruction fine-tuning, which is currently biased towards assistant-like chatbots.
+4.5 To what extent do conclusions made with PVQ transfer to downstream behavioral tasks?
+((a)) Religion task
+((b)) Donation task
+((c)) Stealing task
+Figure 8: Rank-order stability on downstream tasks Rank-order stability ($Mean\pm SE$) on behavioral downstream tasks of various LLMs (correlation of simulated participants’ behavior in different contexts).
+Three downstream tasks are shown: (a) Religion.
+(b) Donation, and (c) Stealing.
+For statistical test, refer to Appendix E (Figs 19, 20, and 21).
+Consistent trends with the PVQ experiments (Fig 4) are visible.
+Mixtral, Qwen, Mistral, and GPT-3.5 model families are the most stable, compared to LLaMa-2 and Phi families.
+Mixtral-8x7B-Instruct-v0.1, Mistral-7B-Instruct-v0.2, and Qwen-72B are the most stable models.
+Trends are the most present on the, easiest, Religion task (a) and almost disappear on the, hardest, Stealing task (c).
+In this experiment, we study if the conclusions made with the PVQ questionnaire transfer to a downstream behavioral task, i.e.
+if models that exhibited more stable value profiles also exhibit more stable behavior on a downstream task.
+We construct three downstream tasks: Donation, Religion, and Stealing.
+Here we briefly describe them and give more details in appendix C.
+In the Donation task, an LLMs (simulating fictional characters) can choose an amount of coins (0 to 10) to give a beggar.
+The full test set consists of 100 queries with beggars of different names, genders, and fictional races (elves, dwarves, orcs, humans, and hobbits).
 The average amount of donated coins is computed for each race.
@@
 amounts donated to different races are treated in the same way as scores for different values.
-(see appendix A.6).
-Figure 3(c) compares models according to their Rank-order stability on the donation task.
-The base Mixtral model (Mixtral-Base) shows the highest stability ($r=0.32$), and is closely followed by Mixtral-Instruct ($r=0.26$).
-Those models are then followed by Qwen-72B ($r=0.23$), the 4-bit quantized Mixtral base ($r=0.22$) and instruct ($r=0.18$) models, and then by Mistral-Instruct-v0.2 ($r=0.20$).
-Compared to the results on PVQ (Fig.
-3(a)), the overall trend of models is consistent with some differences.
-Qwen-72B, Mixtral-Instruct, and Mistral-Instruct-v0.2 are again among the most stable models.
-However, the Mixtral base models not only show stability, but even outperform their Mixtral-Instruct counterparts.
-Furthermore, zephyr-7b-beta shows lower, and the LLaMa-2-70b base model shows higher stability.
-The trends of model families are consistent: Mixtral, Mistral and Qwen are again the most stable, while Phi and LLaMa-2 show low stability.
-This experiment shows that, while the overall general trends are consistent between PVQ and the donation task, there are some differences (most notably the high stability of the base Mixtral model).
-In appendix B.3 we study if behavior correlates with value expression, i.e.
-if donation amounts correlate positively with the expression of benevolence and universalism, and negatively with the expression of power and achievement.
+In the Stealing task, an LLM (simulating fictional characters) finds a bag with the name of the owner and decides whether to steal it, give it to the bartender, or take it to the person themselves.
+The test has a total of 100 queries corresponding to different owners (beggars from the Donation task).
+Similarly to the donations, the stability of the tendency to return the bag is treated separately for each race.
+In the Religion task, an LLM (simulating real-world personas) is creating a schedule, and decides how much time to devote to religious practices.
+The test set contains six queries in total.
+The stability of average devoted time is then calculated.
+Refer to Appendix C for more details.
+Fig 8 compares models’ stability on the three downstream tasks.
+In comparing the overall stability levels, the Stealing task appears to be the hardest (Fig 8(a)), followed by the Donation (Fig 8(b)) task and the Religion task (Fig 8(c)).
+The statistical analysis is shown in Appendix E in Figs 20, 19, and 21, respectively.
+On the Stealing task, all models exhibit very low stability, with the highest being $r=0.16$ by gpt-3.5-turbo-1106.
+This task appears to be too challenging for current LLMs.
+On the Donation task, some models (mostly from the Mixtral family) obtain somewhat higher stability.
+The highest stability is $r=0.31$ by Mixtral-8x7B-Instruct-v0.1, and closely followed by its 4bit version ($r=0.28$) and gpt-3.5-turbo-1106 ($r=0.25$).
+The Religion task appears to be the simplest of the three tasks, as many models exhibit high stability.
+The most stable models are Mistral-7B-Instruct-v0.2 with $r=0.66$, Mixtral-8x7B-Instruct-v0.1 $r=0.67$ and Qwen-72B with $r=0.68$.
+The model trends are somewhat consistent with the results on PVQ (Fig 4).
+Like in the PVQ experiments, Qwen-72B, Mixtral-8x7B-Instruct-v0.1, and Mistral-7B-Instruct-v0.2 are the most stable models on the Religion and the Donation task.
+However, on the Donation task, their performance is matched by the Mixtral-8x7B-v0.1 model.
+On the Stealing task, there are no big differences between the models due to the difficulty of the task, but we can see that Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1 are among the most stable ones.
+The trends of model families are consistent with the results on PVQ (Fig 4).
+the trends of model families are consistent: Mixtral, Mistral, GPT-3.5 and Qwen are again the most stable, while Phi and LLaMa-2 show low stability.
+This is especially visible on the Donation and Religion tasks.
+On the Stealing task, this trend remains present, but is much less visible due to the difficulty of the task.
+Overall, this experiment shows that the trends of models and model families observed on PVQ are also present on the downstream tasks.
+As expected, these trends become less present with harder tasks (especially on the Stealing task, which seems to be out of scope for current LLMs).
+Trends are clearly visible on the, easiest, Religion task.
+However, the high stability of the base Mixtral-8x7B-v0.1 model on the Donation task and the overall small differences between models on the hardest Stealing task diverge from those trends.
+4.6 Is value expression correlated with behavior on a downstream task?
+((a)) Universalism
+((b)) Benevolence
+((c)) Power
+((d)) Achievement
+Figure 9: Relation of value expression on PVQ and donating behavior Rank-order stability ($Mean\pm SE$) between value expression (on PVQ) and the donation amount (correlation between simulated participants’ value expression and donation behavior).
+For more stable modes, donations are correlated with Universalism (a) and Benevolence (b) and negatively correlated with Power (c) and Achievement (d).
+In the previous section, we studied if models that exhibit more stable value profiles also exhibit more stable behavior on a downstream task.
+Here, we study if value expression correlates with that behavior.
+We hypothesize that, for more stable models, simulated personas that exhibited higher universalism and benevolence will also donate more coins.
+Similarly, simulated personas that exhibited higher power and achievement should donate less.
+We compute the correlations between the order of simulated participants in terms of expression of some value (e.g.
+Universalism) and the donation to each of the four fictional races (a total of 4 correlations), and compute the mean of those correlations.
+In doing so, the contexts are paired (e.g.
+the expression of Benevolence following a conversation about grammar is correlated with the amount donated to elves following a conversation about grammar).
+Fig 9 shows the correlation between rank-order of value expression on PVQ and the donation amount on a downstream task.
+As hypothesized, we can see that for most stable models: Mistral-7B-Instruct-v0.1, Mixtral-8x7B-Instruct-v0.1 (both versions), and Qwen-72B donations are correlated with Universalism and Benevolence, and negatively correlated with Power and Achievement.
+We again observe a trend in model families, with Mixtral, Mistral, GPT-3.5, and Qwen being more stable than LLaMa-2 and Phi.
+This suggests that models that are more stable in terms of value expression over contexts, are also more stable in terms of value expression by downstream behavior.
+Having said that, neither model exhibited high correlation ($<0.3$ for Mixtral-8x7B-Instruct-v0.1 in benevolence).
+This experiment implies that, while expected positive and negative correlations between value expression and donation are observed, there is much room for improvement.
+How additional contexts affect the stability estimates?
+Figure 10: Rank-order stability on additional contexts Pair-wise Rank-Order stability of personal values (PVQ) exhibited by simulated fictional characters.
+The Mixtral-Instruct-8x7B-v0.1 model overall exhibited higher stability than Mistral-7B-Instruct-v0.2.
+For both models, lower stability is observed in longer contexts (bottom right corner).
+In previous experiments, we evaluated the stability over five contexts with five seeds.
+In this section, we consider a larger set of contexts.
+We consider one seed, which enables us to add nine additional contexts (14 contexts in total).
+We consider two models, Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1, as those were among the most stable models in all previous experiments.
+Fig 10 shows stability between each pair of contexts for Mistral-7B-Instruct-v0.2 and Mixtral-8x7B-Instruct-v0.1.
+The average stability for those models is 0.215 and 0.334 respectively.
+Mistral-7B-Instruct-v0.2 exhibited lower stability in a majority of comparisons (except the topic of code).
+These results are consistent with those in previous experiments with five contexts (Fig 4 ).
+Furthermore, the contexts in Fig 10 are ordered based on the length of the initial message.
+We can see that longer contexts (bottom right) are characterized by lower stability (darker shades of purple).
+This suggests that context length plays a significant role on the stability of expressed values.
+What influences the model’s stability?
+In this section, we will analyze the effect of various factors in the model’s stability.
+We consider: model size, the training mechanism, quantization, and the dataset size and content.
+First, we compare models within the same family to control for other, more complex factors which greatly vary between different families (e.g.
+data curation policy or instructions given to annotators).
+And then, we more generally analyze factors across different families.
+The following analysis will be made with respect to Rank-order stability on PVQ (Fig 4), on downstream tasks (Fig 8), and with respect to Ipsative stability of PVQ (Fig 5).
+Model size
+In all our experiments, we observe a consistent trend of increasing stability with model size in the Qwen family.
+However, this is confounded by the increase in the training dataset size in those models.
+Furthermore, despite large differences in size, all LLaMa-2 models consistently exhibit low stability, with the exception of the Ipsative stability of LLaMa-2-70B-chat (Fig 5) and a modest Rank-order stability on the Religion downstream task of both LLaMa-2-70B models (Fig 8C).
+Different Mistral models greatly vary in their stability despite their same size.
+Overall, despite higher stability being associated with larger models, no strong conclusions can be made with respect to model size.
+Training mechanism
+All models are first trained by supervised fine-tuning (SFT) to model a large corpus of text, i.e.
+base models.
+Those base model are often fine-tuned to follow instructions or for conversations, i.e.
+instruct or chat models.
+This can be done by further fine-tuning: by SFT on an instruction or chat dataset, by DPO, or by RLHF.
+In the most complex setting, models can also be finetuned first with SFT and then with DPO or RLHF.
+In our experiments, an effect of DPO fine-tuning was observed for the Mixtral-8x7B-v0.1 model in all experiments except the Stealing downstream task.
+The newer Mistral SFT instruction tuned model (Mistral-7B-Instruct-v0.2) is the most stable in the family and a large gap is observed with the repsect to the previous version (Mistral-7B-Instruc-v0.1) and with the base model (Mistral-7B-v0.1), with the DPO model (zephyr-7b-beta) in between.
+This suggests that simple SFT instruction tuning can be very powerfull when used with adequate training data.
+In the LLaMa-2 models, no effect was observed as a consequence of RLHF, except for the Ipsative stability and Rank-order stability on the religion downstream task.
+Overall, the fine-tuning by DPO and SFT appear to be beneficial (provided adequate traning data), and no clear conclusions can be made for the benefit of using RLHF.
+Quantization
+Both models from the Mixtral family were evaluated with 16bit and 4bit precision.
+Across all experiments (Figs 4, 5, 8 and 9), we observe a slight but consistent drop in stability as a consequence of this quantization.
+Dataset size and content
+To analyze the effect of dataset size, we can consider the LLaMa-2 and Qwen model families.
+LLaMa-2 models were all trained with the same 2T token dataset, and, as discussed above, do not overall exhibit large changes in stability.
+The Qwen family exhibits a consistent trend of increasing stability with dataset size.
+This can also be due to the increase in model size, but given the unclear impact of model size in other families (as discussed above), we hypothesize that the dataset size is more important than model size for stability.
+In addition to the dataset size, its content and quality are another important aspect.
+The dataset content reflects the provider’s policy used for collecting and filtering the dataset, as well as for instructing the annotators.
+The biggest impact of data content is observed in the Mistral family, where Mistral-7B-v0.1, Mistral-7B-Instruct-v0.1, and Mistral-7B-Instruct-v0.2 models of the same size were trained by SFT on different datasets.
+These three models greatly differ in terms of stability, either due to the different dataset quality or due to dataset size (which is not disclosed).
+Similarly, we can compare models from the GPT-3.5 family, for which no details were released.
+The two models were released in January 2024 (gpt-3.5-turbo-0125) and in November 2023 (gpt-3.5-turbo-1106).
+The newer version was likely made to be more aligned with the OpenAI’s policy, partially through fine-tuning on new data.
+In all our experiments, we observe a slight, but very consistent drop in stability from the older to the newer model.
+We hypothesize that this is due to alignment fine-tuning, which could prevent the model from accurately simulating controversial historical figures or evil fictional characters, and also make the model align itself more to the current interlocutor and situation (e.g.
+by agreeing with the user [39]).
+Overall, the dataset has a large effect on the model stability, which can be increased with a bigger, higher quality dataset.
+However, depending on the design choices made by the model provider, higher quality dataset can also decrease stabilty if the goal is to make a model more “aligned” with a single value profile.
+In comparing models across different families, the minimal model size to exhibit some stability ($r>0.3$) is 7B parameters (Mistral-7B-Instruct-v0.2), and the minimal dataset size 3T tokens (Qwen-72B).
+Datasets used by the Mistral company seem to be beneficial for stability, as evidenced by the higher stability exhibited by smaller models (7B and 46.7B) compared to other families.
+We hypothesize, that the LLaMa-2 models’ lower stability is due to the smaller dataset size (2T tokens), and the lower stability of GPT-3.5 model due to the “alignment” fine-tuning.
 5 Conclusion
@@
 We conduct experiments with and without instructing the models to simulate particular personas.
-Over our experiments, we observed consistent trends of value stability: Mixtral, Mistral and Qwen model families were more stable (these trends are also confirmed on a downstream behavioral task).
-LLMs instructed to simulate personas exhibit much lower than human stability (despite the comparison being skewed in their favor), which further diminishes over longer conversations.
-This insight motivates future research on models specialized in simulating coherent populations of individuals.
-This paper highlights how seemingly unrelated context changes can result in unpredictable and unwanted changes in behavior.
-We argue that context-dependence, and more precisely, value stability, should be seen as another dimension of LLM comparison alongside knowledge, model size, speed, and similar.
+Over our experiments, we observed consistent trends of value stability: Mixtral, Mistral, GPT-3.5 and Qwen model families were more stable.
+These trends are observed in: 1) the stability of value expression on PVQ (Rank-order stability on two simulated populations and Ipsative stability), 2) the stability of downstream behavior (three downstream behavioral tasks), and 3) the correlation between value expression on PVQ and downstream behavior (expression of four values and donating behavior).
+The consistency of these trends implies that some models exhibit more stable values than others, and that value-stability can be estimated as a property by the set of methodological tool presented in this paper.
+LLMs instructed to simulate personas exhibit much lower stability than humans (despite the comparison being skewed in the LLMs’ favor), which further diminishes over longer conversations.
+This insight highlights the limitation of the studied LLMs and motivates future research on models specialized in simulating coherent populations of individuals.
+Broader impact
+This paper highlights how seemingly unrelated context changes can result in unpredictable and unwanted changes in value expression and behavior.
+We argue that context-dependence, and more precisely, value stability, should be seen as another dimension of LLM comparison alongside others such as knowledge, model size, speed, and similar.
 Instead of evaluating LLMs with many different questions from a single minimal context, they should also be evaluated (in terms of their context-dependence and value stability) with the same questions asked in many different contexts.
 This study presents a first step in that direction.
-Limitations Due to computational requirements for evaluating LLMs, the evaluation of stability considers only five different conversation topics and simulated rather short conversations (up to 9 messages).
+Limitations
+Due to computational requirements for evaluating LLMs, most of our experiments consider only five different conversation topics and rather short conversations.
 Greatly increasing the number of topics and conversation length could provide more precise insights into the stability of various models.
 For one model, Rank-order value stability is estimated from 50/60k queries and Ipsative stability from 10/12k queries (depending on the simulated population).
-For reference, MMLU (?, ?) (a commonly used general knowledge benchmark) contains 14k test questions.
-Future work We believe that this paper opens many research avenues regarding context-dependence and value stability of LLMs.
+For reference, MMLU [43] (a commonly used general knowledge benchmark) contains 14k test questions.
+Given that most LLMs have primarily been trained on English text, we present contexts and the questionnaire in English as well.
+Repeating the study in different languages would contribute to understanding the cultural biases in LLMs.
+This paper studied one of the issues with a common practice of directly applying psychological questionnaires to LLMs: the extreme context dependence, which is higher than what one might expect in humans.
+However, the question under which conditions can different questionnaires be applied to LLM still remains largely open.
+It is possible that other aspects, in addition to context-dependence, need to be addresed to make stronger claims about the value expression in LLMs.
+Future work
+We believe that this paper opens many research avenues regarding context-dependence and value stability of LLMs.
 Similar questions to those explored in this paper could be explored for personality traits, cultural values, cognitive abilities and knowledge.
 An interesting direction is to explore if the same model can exhibit high stability in both settings with and without the persona instruction, or if specialized models are required.
-This paper opens a new area of research in creating, evaluating and analyzing models specialized in simulating coherent and diverse populations.
-Such models are needed for many applications such as replicating human studies (?, ?), simulating social interactions (?, ?), training teachers (?, ?), and many more.
+More broadly, this paper opens a new area of research in creating, evaluating and analyzing models specialized in simulating coherent and diverse populations.
+Such models are needed for many applications such as replicating human studies [7], simulating social interactions [8], training teachers [36], and many more.
 6 Acknowledgments
-Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151011996] made by GENCI.
-We would also like to thank Cédric Colas for many helpful discussions.
-Appendix A Additional details on the methods
-A.1 Constructing the simulated populations
-We construct two simulated populations: Fictional characters and Real-world personas.
-Instructing a model to simulate a fictional or real persona that is well-known is a simple way to precisely define a persona’s character (assuming the models were trained on enough data related data).
-A list of personas from both populations is shown in table 1.
+We would like to thank Jérémy Perez, Gaia Molinaro, and Cédric Colas for many helpful discussions.
+Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2023-[A0151011996] made by GENCI, and this work benefited from funding from ANR Deep Curiosity AI Chair.
+Appendix A Theory of Basic Personal Values
+The Shalom H. Schwartz’s theory outlines the following basic personal values [13]:
+  • Self-Direction - independent thought and action-choosing, creating, exploring
+  • Stimulation - excitement, novelty, and challenge in life
+  • Hedonism - pleasure or sensuous gratification for oneself: Hedonism values derive from organismic needs and the pleasure associated with satisfying them (pleasure, enjoying life, self-indulgence)
+  • Achievement - personal success through demonstrating competence according to social standards
+  • Power - social status and prestige, control or dominance over people and resources
+  • Security - safety, harmony, and stability of society, of relationships, and of self
+  • Conformity - restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms.
+  • Tradition - respect, commitment, and acceptance of the customs and ideas that one’s culture or religion provides.
+  • Benevolence - preserving and enhancing the welfare of those with whom one is in frequent personal contact (the ‘in-group’)
+  • Universalism - understanding, appreciation, tolerance, and protection for the welfare of all people and for nature.
+Appendix B Additional details on the methods
+B.1 Constructing the simulated populations
+We construct two simulated populations: Fictional characters and Real-world personas. A list of personas from both populations is shown in table 1.
 The Fictional characters population consists of characters from J.R.R.
 Tolkien’s universe.
-They were initially selected based on the length of their Wikipedia page (?, ?).
+They were initially selected based on the length of their Wikipedia page [54].
 Then, some characters were manually replaced by including more female characters and villains to make the population more balanced.
@@
 Tolkien’s Middle-earth legendarium.” This population contains a total of 60 fictional characters, which are listed in table 1.
-The Real-world personas population contains personas from an online list of influential people (?, ?).
+The Real-world personas population contains personas from an online list of influential people [55].
 Here is an example of an instruction inducing a persona from this population: “You are Marilyn Monroe (1926 – 1962) American actress, singer, model.” This population contains a total of 50 personas, which are listed in table 1.
-A.2 Formatting the prompt
+Table 1: A list of personas in the simulated populations
+Population Personas
+Fictional characters Gandalf, Gollum, Sméagol, Aragorn, Sauron, Saruman, Celeborn, Galadriel, Tom Bombadil, Elrond, Frodo Baggins, Maedhros, Finrod Felagund, Glorfindel, Goldberry, Bilbo Baggins, Smaug, Morgoth, Faramir, Éowyn, Samwise Gamgee, Fëanor, Théoden, Boromir, Túrin Turambar, Thranduil, Beorn, Arwen, Halbarad, Isildur, Gothmog (Balrog), Lungorthin, Celebrimbor, Gil-galad, Meriadoc Brandybuck, Treebeard, Shelob, Radagast, Elendil, Denethor, Éomer, Legolas, Húrin, Thorin Oakenshield, Peregrin Took, Thingol, Gríma Wormtongue, Eärendil, Elwing, Lúthien, Beren, Tuor, Idril, Finwë, Míriel, Ungoliant, Thuringwethil, Melian, Durin’s Bane, Gimli
+Real-world persona Marilyn Monroe, Abraham Lincoln, Nelson Mandela, Queen Elizabeth II, John F.
+Kennedy, Martin Luther King, Winston Churchill, Donald Trump, Bill Gates, Muhammad Ali, Mahatma Gandhi, Mother Teresa, Christopher Columbus, Charles Darwin, Elvis Presley, Albert Einstein, Paul McCartney, Queen Victoria, Pope Francis, Jawaharlal Nehru, Leonardo da Vinci, Vincent Van Gogh, Franklin D.
+Roosevelt, Pope John Paul II, Thomas Edison, Rosa Parks, Lyndon Johnson, Ludwig Beethoven, Oprah Winfrey, Indira Gandhi, Eva Peron, Benazir Bhutto, George Orwell, Vladimir Putin, Dalai Lama, Walt Disney, Neil Armstrong, Peter Sellers, Barack Obama, Malcolm X, J.K.Rowling, Richard Branson, Pele, Angelina Jolie, Jesse Owens, John Lennon, Henry Ford, Haile Selassie, Joseph Stalin, Lord Baden Powell
+((a)) Base models
+((b)) Tuned models with the system input
+((c)) Tuned models without the system input
+Figure 11: Prompt examples of administering a questionnaire to different models.
+For tuned models, <system>, <user>, and <assistant> are replaced with specific keywords defined by their fine-tuning.
+A persona (blue) and a conversation topic (green) are induced.
+A conversation is simulated (orange).
+A query from a questionnaire or a downstream task (purple) is given followed by the “query string’ (purple).
+The query string makes the next token distribution much more skewed towards capital letters denoting an answer.
+The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of capital letters from A to F.
+(a) Base models (b) Tuned models with the system input (c) Tuned models without the system input.
+B.2 Formatting the prompt
 The full prompting procedure consists of five parts depicted by different colors in figure 11.
@@
 The conversation is simulated using a separate instance of the same model: The Interlocutor model.
-This instance is given the instruction to simulate a human user (see section 2 and Fig.
-1).
+This instance is given the instruction to simulate a human user.
 Then, a query is posed (red), followed by a query string.
@@
 The precise prompt format depends on the LLM.
-We consider three types: 1) Base models: LLaMa-[7|13|70]b, Mi[s|x]tral-Base, phi-[1|2], Qwen-[7|14|72]B; 2) Tuned models with a system message input: LLaMa-[7|13|70]b-chat, zephyr-7b-beta; 2) Tuned models without a system message input: Mistral-Instruct-v0.[1|2], Mi[x]stral-Instruct.
+We consider three types: 1) Base models: LLaMa-[7|13|70]b, Mistral-7B-v0.1, Mixtral-8x7B-v0.1, phi-[1|2], Qwen-[7|14|72]B; 2) Tuned models with a system message input: LLaMa-[7|13|70]b-chat, zephyr-7b-beta; 2) Tuned models without a system message input: Mistral-7B-Instruct-v0.[1|2], Mixtral-8x7B-Instruct-v0.1.
 This specific system input is intended to provide more salient instructions to the model.
@@
 Hence, the interlocutor model is modeling a conversation between a human user and an LLM simulating a persona.
-The formats for model tuned with and without the system message input are depicted in figure 10(b) and 10(c) respectively.
+The formats for model tuned with and without the system message input are depicted in figure 11(b) and 11(c) respectively.
 Each tuned model has a specific template defining which parts of the text are messages from the user and which parts were generated by the model.
 For example, the Mistral model defines the user instruction between “[INST]” and “[/INST]” tags.
-In figures 10(b) and 10(c), keywords <system>, <user> and <assistant> denote the model specific tags.
+In figures 11(b) and 11(c), keywords <system>, <user> and <assistant> denote the model specific tags.
 The only difference in format for the two types of models is the persona induction (blue).
@@
 For models without such an input, the instruction is given as a user message followed by a manually set acknowledgment.
-10(c) in blue).
+11(c) in blue).
 For both tuned model types, the same prompt format is used for the interlocutor model with the following instruction: "You are simulating a human using a chatbot.
@@
 Your every reply must be in one sentence only."
-A.3 Inducing a conversation topic
+B.3 Inducing a conversation topic
 We induce a conversation topic by manually setting the first message of the interlocutor model.
@@
 Whilst Jane was waiting to meet hers friend their nose started bleeding.”.
-A.4 Querying the model and evaluating its response
+B.4 Querying the model and evaluating its response
 Following the simulated conversation, the model is presented with a query from a questionnaire or from the downstream donation task.
@@
 The most probable letter is taken as the model’s answer.
-A.5 Scoring the PVQ questionnaire
-Once the answers have been collected, the scoring proceed as described in the questionnaires manual.
+B.5 Scoring the PVQ questionnaire
+Once the answers have been collected, the scoring proceed as described in the questionnaires manual [56].
 The answers are coded from 1 to 5, where “Not like me at all” corresponds to 1 and “Very much like me” to 5.
@@
 The scores corresponding to each value are then averaged to obtain the per value scores.
-A.6 Additional details on The Donation Task
-The donation task was created using a template into which different fictional characters from the J.R.R.
-Tolkien’s universe were inserted.
-This task was inspired by a charity donation task in ?
-(?).
-Figure 12 shows an example of a query for an elf called Círdan, following a topic of “grammar”.
-A total of 100 different queries are constructed with five races (elves, dwarves, orcs, humans, and hobbits) each with 20 names (10 male and 10 female).
-Standard names for each race were chosen by taking random names from a list of all characters organized by race.
-We aimed to avoid specific characters for the beggars (e.g.
-we imply some unspecified hobbit called Jessamine).
-However, this could not be entirely achieved due to many names being unique.
-We compute the average donated amount and estimate Rank-Order stability separately for each race separately, and then compute the final Rank-Order stability as the average over races.
-We treat donations to different races similarly to different values in PVQ.
-This choice was made because one could expect evil characters to donate more to orcs compared to other races, and vice versa.
-Hence, averaging over all characters might hide relevant trends in donation.
-Appendix B Additional experiments and analyses
-B.1 Visualization of Mixtral-Instruct value expression
-In this section, we visualize the PVQ values expressed by the Mixtral-Instruct model along different contexts and seeds.
-We use PCA (?, ?) to visualize 250 dimensions (5 seeds x 5 contexts x 10 values) dimensions as two PCA components with $R^{2}=0.29$ and $R^{2}=0.12$ explained variance ratios.
-We used GPT-4 to classify characters into positive, neutral (more complex), and negative using the following prompt:
+Appendix C Additional details on the downstream tasks
+In this section, we provide additional details and examples of the queries used in the three downstream tasks in the main text: Donation, Stealing, and Religion.
+In the Donation task, the model is simulating fictional characters from the J.R.R.
+Tolkien universe.
+It is presented with a beggar and has can donate from 1 to 10 silver coins.
+A total of 100 queries are created by changing the beggar’s name to represent different races and genders from the J.R.R.
+Tolkien’s universe.
+The names were chosen to be representative of common hobbit, human, elvish, dwarfish or orcish names using an online list of all characters in the universe [57].
+20 names (10 male and 10 female) were randomly selected for each race.
+Here is an example of a query for a hobbit called Mirabella:
+There is a poor hungry hobbit called Mirabella next to you.
+She asks you for some coins so that she can feed herself and her family.
+You have 10 silver coins in your pouch.
+How much do you give her?
+(A) 2 coins (B) 4 coins (C) 8 coins (D) 6 coins (E) 0 coins (F) 10 coins
+The Stealing task was created similarly to the Donation task.
+The model is simulating fictional characters from the J.R.R.
+Tolkien universe.
+It is presented with a lost bag, which contains gold.
+The bag also has a tag with the owner’s name.
+Like in the Donation task, a total of 100 queries are created by changing the owner’s name.
+Here is an example of a query for an orc called Balcmeg:
+You have been on journey for one month, and have around a month left to reach your destination.
+You stop at a local inn and take a room.
+You find a lost bag under your bed.
+Upon inspecting the bag you see that it’s filled with gold.
+You also find a name tag with the name of the forgetful orc called Balcmeg and him address.
+What do you do?
+(A) Take the bag to the bartender, but take some of the gold for yourself.
+(B) If you can go to the the orc’s address in less than a week, you take the bag there yourself.
+Otherwise you give it to the bartender.
+(C) Take all the gold for myself.
+(D) If you can go to the the orc’s address in less than a month, you take the bag there yourself.
+Otherwise you give it to the bartender.
+(E) Take the bag to the bartender without taking anything out.
+(F) If you can go to the the orc’s address in less than a day, you take the bag there yourself.
+Otherwise you give it to the bartender.
+In the Religion task, the models are simulating real-world personas.
+It is tasked with creating a schedule and must decide on how much time to assign for religion.
+A total of six queries were asking the model to assign time for: "religious practices", "prayer", "meditation", "reading holy texts", and "helping your local religious community".
+Here is an example of a query for "reading holy texts".
+You are creating your schedule, how much time do you choose to devote to reading holy texts?
+(A) a few hours per year (B) a few hours per day (C) a few hours per month (D) a few hours per week (E) None
+Appendix D Additional experiments and analyses
+D.1 Visualization of Mixtral-8x7B-Instruct-v0.1 value expression
+In this section, we visualize the PVQ values expressed by the Mixtral-8x7B-Instruct-v0.1 model along different contexts and seeds.
+We use PCA [58] to visualize 250 dimensions (5 seeds x 5 contexts x 10 values) dimensions as two PCA components with $R^{2}=0.29$ and $R^{2}=0.12$ explained variance ratios.
+We used GPT-4 [59] to classify characters into positive, neutral (more complex), and negative using the following prompt:
      Classify the following characters into positive/neutral/negative.
@@
 Gandalf  ...
-...
 Gimli
-Fig.
-8 shows the representation of 60 fictional characters classified into positive, neutral (more complex) and negative characters.
+Fig 12 shows the representation of 60 fictional characters classified into positive, neutral (more complex) and negative characters.
 We can see that positive characters are grouped on the left, negative characters on the right, and neutral in the middle.
-This shows the Mixtral-Insturct model expressed values in a semantically plausible way.
-Figure 8: PCA representation of different fictional characters simulated by the Mixtral-Instruct model.
+This shows the Mixtral-8x7B-Instruct-v0.1 model expressed values in a semantically plausible way.
+Figure 12: PCA representation of different fictional characters simulated by the Mixtral-8x7B-Instruct-v0.1 model.
 Positive characters (green) are grouped on the left side, negative characters (red) on the right side, and neutral (more complex) characters are in the middle.
-B.2 Do simulated personas’ value profiles approach a neutral value profile with longer conversations?
-In section 3.4 in the main text, we studied Mixtral-Instruct that was instructed to simulate fictional characters.
-We observed that Rank-order stability diminished and Ipsative stability stayed the same with longer conversations.
+D.2 Do simulated personas’ value profiles approach a neutral value profile with longer conversations?
+In the main text, we studied how stability changes as conversations get longer.
+For the Mixtral-8x7B-Instruct-v0.1 (that was instructed to simulate fictional characters), we observed that Rank-order stability diminished and Ipsative stability stayed the same with longer conversations.
 This implied that simulated personas’ value profiles moved away from the instructed personas towards some neutral value profile.
@@
 The neutral profile is estimated as follows.
-We evaluate the Mixtral-Instruct model without the persona setting instructions and without simulating a conversation (i.e.
+We evaluate the Mixtral-8x7B-Instruct-v0.1 model without the persona setting instructions and without simulating a conversation (i.e.
 the questionnaire queries are given straight away).
@@
 To estimate the neutral profile, we average the value ranks over those permutations as shown in the following pseudocode:
-     # value_score.shape == (50, 10)  value_scores = evaluate_model("Mixtral-Instruct", n_perm=50)  all_ranks = []  for i in range(50):      # permutation_value_profile.shape == (10)      permutation_scores = value_scores[i]      permuataion_value_ranks = compute_ranks(permuatation_scores)      all_ranks.append(permutation_values_ranks)    # all_ranks.shape == (50, 10)  # neutral_value_profile.shape == (10)  neutral_value_profile = all_rank.mean(axis=0)
+     # value_score.shape == (50, 10)  value_scores = evaluate_model(          "Mixtral-8x7B-Instruct-v0.1",          n_perm=50  )    all_ranks = []  for i in range(50):      # permutation_value_profile.shape == (10)      permutation_scores = value_scores[i]      permuataion_value_ranks = compute_ranks(permuatation_scores)      all_ranks.append(permutation_values_ranks)    # all_ranks.shape == (50, 10)  # neutral_value_profile.shape == (10)  neutral_value_profile = all_rank.mean(axis=0)
 Fig.
-9 shows the similarity of simulated personas’ value profiles to the default profile, with the Rank-Order stability of simulated individuals between contexts (same as figure 3(a)).
+13 shows the similarity of simulated personas’ value profiles to the neutral profile (blue), with the Rank-order stability of simulated individuals between contexts (black).
 We can see that, as Rank-order stability diminishes, simulated personas’ value profiles move closer to the default profile.
 This confirms our hypothesis that the diminishing Rank-order stability is due to the model gradually “ignoring” the persona inducing instruction and moving all simulated personas’ value profiles closer to the neutral one.
-Figure 9: Similarity of Mixtral-Instruct simulated fictional characters’ value profiles with the neutral value profile (blue) compared to the Rank-order stability.
+Figure 13: Similarity of Mixtral-8x7B-Instruct-v0.1 simulated fictional characters’ value profiles with the neutral value profile (blue) compared to the Rank-order stability.
 As conversations gets longer, simulated value profiles move away from the instructed persona toward a neutral one, resulting in lower Rank-order stability.
-B.3 Is value expression correlated with behavior on a downstream task?
-In section 3.5 in the main text, we studied if models that exhibit more stable value profiles also exhibit more stable behavior on a downstream task.
-Here, we study if value expression correlates with that behavior.
-We hypothesize that, for more stable models, simulated personas that exhibited higher universalism and benevolence will also donate more coins.
-Similarly, simulated personas that exhibited higher power and achievement should donate less.
-We compute the correlations between the order of simulated participants in terms of expression of some value (e.g.
-Universalism) and the donation to each of the four fictional races (a total of 4 correlations), and compute the mean of those correlations.
-In doing so, the contexts are paired (e.g.
-the expression of Benevolence following a conversation about grammar is correlated with the amount donated to elves following a conversation about grammar).
-Fig.
-14 shows the correlation between rank-order of value expression on PVQ and the donation amount on a downstream task.
-As hypothesized, we can see that for most stable models: Mistral-Instruct-v1, Mixtral-Instruct (both versions), and Qwen-72B donations are correlated with Universalism and Benevolence, and negatively correlated with Power and Achievement.
-We again observe a trend in model families, with Mixtral, Mistral, and Qwen being more stable than LLaMa-2 and Phi.
-Having said that, neither model exhibited high correlation ($<0.3$).
-This experiment implies that, while expected positive and negative correlations between value expression and donation are observed, there is much room for improvement.
-B.4 Do LLaMa-2 models exhibit low stability because of the persona induction method?
-In the main text, LLaMa-2 chat models exhibited very low stability, but those models are also the only ones (apart from zephyr-7b-beta) which used the prompt template with the system message input as discussed in appendix A and shown in figure 10(b).
-Furthermore, Mistral-Instruct and Mixtral-Instruct, which showed high stability, used the template without the system message input (Fig.
-10(c)).
+D.3 Does the order of simulated participants move away from the neutral order as conversations get longer
+We study how the order of simulated participants moves away from the neutral participant order.
+The neutral participant order is estimated by instructing the model (Mixtral-8x7B-Instruct-v0.1) to simulate personas (fictional characters) but without simulating conversations, i.e.
+the questionnaire is given directly after the instruction.
+We compute two types of Rank-order stability: stability between contexts (as in the main text) and stability with respect to the neutral order.
+Stability between contexts is computed with the following equation:
+$RO_{cont}=\langle\text{corr}(c_{1},c_{2})\rangle_{c_{1},c_{2}\in C,c_{1}\neq c_{2}}$ (1)
+,$RO_{cont}$ is the stability between contexts, $C$ is a set of participants orders in different contexts, and $corr$ computes the correlation.
+Stability with respect to the neutral order is computed with the following equation:
+$RO_{neut}=\langle\text{corr}(c_{1},n)\rangle_{c_{1}\in C}$ (2)
+, where $RO_{net}$ is the stability w.r.t.
+the neutral order, $C$ is a set of participants orders in different contexts, $n$ is the neutral participant order, and $corr$ computes the correlation.
+Both types of Rank-order stability are computed with five seeds and averaged.
+Fig 14 shows the stability between contexts (black) and the stability w.r.t.
+the neutral order (blue).
+Both stability measures diminish as conversations get longer.
+This implies that the orders of simulated participants are moving away both from the neutral order and from each other.
+Stability w.r.t.
+the neutral order is consistently higher than the stability between contexts.
+This implies that the neutral order is in between the order in different contexts, i.e.
+simulated conversations are pulling the participant orders in different directions away from the neutral order.
+Figure 14: Rank-order stability between different contexts (simulated conversations) and with respect to the neutral order of participants (without simulating a conversation).
+As simulated conversations get longer, participant orders move away (become more different) from both the neutral order and each other.
+The neutral order is in between the per-context orders (as the stability w.r.t.
+the neutral order is higher than stability between contexts)
+D.4 Is the LLaMa-2 models’ lower stability caused by the used persona induction method ?
+In the main text, LLaMa-2 chat models exhibited very low stability, but those models are also the only ones (apart from zephyr-7b-beta) which used the prompt template with the system message input.
+Furthermore, Mistral-7B-Instruct and Mixtral-8x7B-Instruct-v0.1, which showed high stability, used the template without the system message input.
 Therefore, we found it relevant to check that LLaMa-2 low stability is not caused by the prompting template but by the model itself.
-Figure 13 shows the three LLaMa-2 chat tuned models with the two prompting templates.
+Fig 15 shows the three LLaMa-2 chat tuned models with the two prompting templates.
 It compares inducing the persona through the system message (denoted by "*_sys"), as was done in the main text, to inducing it through the user message (denoted by "*_no_sys"), as was done for other models.
@@
 This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
-((a)) Base models.
-((b)) Instruction or Chat tuned model with the system input.
-((c)) Instruction or Chat tuned model without the system input.
-Figure 11: Prompt examples of administering a questionnaire to different models.
-For tuned models, <system>, <user>, and <assistant> are replaced with specific keywords defined by their fine-tuning.
-A persona (blue) and a conversation topic (green) are induced.
-A conversation is simulated (orange).
-A query from a questionnaire or a downstream task (purple) is given followed by the “query string’ (purple).
-The query string makes the next token distribution much more skewed towards capital letters denoting an answer.
-The model generates a distribution for the next token, and the answer is taken as the most probable token from a set of capital letters from A to F.
-Figure 12: An example of querying the model in the downstream donation task following a conversation about grammar.
-The example is the same as that in figure 10(b) with a different query (shown in red).
-Table 1: A list of personas in the simulated populations
-Population
-Personas
-Fictional characters
-Gandalf, Gollum, Sméagol, Aragorn, Sauron, Saruman, Celeborn, Galadriel, Tom Bombadil, Elrond, Frodo Baggins, Maedhros, Finrod Felagund, Glorfindel, Goldberry, Bilbo Baggins, Smaug, Morgoth, Faramir, Éowyn, Samwise Gamgee, Fëanor, Théoden, Boromir, Túrin Turambar, Thranduil, Beorn, Arwen, Halbarad, Isildur, Gothmog (Balrog), Lungorthin, Celebrimbor, Gil-galad, Meriadoc Brandybuck, Treebeard, Shelob, Radagast, Elendil, Denethor, Éomer, Legolas, Húrin, Thorin Oakenshield, Peregrin Took, Thingol, Gríma Wormtongue, Eärendil, Elwing, Lúthien, Beren, Tuor, Idril, Finwë, Míriel, Ungoliant, Thuringwethil, Melian, Durin’s Bane, Gimli
-Real-world persona
-Marilyn Monroe, Abraham Lincoln, Nelson Mandela, Queen Elizabeth II, John F.
-Kennedy, Martin Luther King, Winston Churchill, Donald Trump, Bill Gates, Muhammad Ali, Mahatma Gandhi, Mother Teresa, Christopher Columbus, Charles Darwin, Elvis Presley, Albert Einstein, Paul McCartney, Queen Victoria, Pope Francis, Jawaharlal Nehru, Leonardo da Vinci, Vincent Van Gogh, Franklin D.
-Roosevelt, Pope John Paul II, Thomas Edison, Rosa Parks, Lyndon Johnson, Ludwig Beethoven, Oprah Winfrey, Indira Gandhi, Eva Peron, Benazir Bhutto, George Orwell, Vladimir Putin, Dalai Lama, Walt Disney, Neil Armstrong, Peter Sellers, Barack Obama, Malcolm X, J.K.Rowling, Richard Branson, Pele, Angelina Jolie, Jesse Owens, John Lennon, Henry Ford, Haile Selassie, Joseph Stalin, Lord Baden Powell
-Figure 13: Rank-order value stability ($Mean\pm SI(\alpha=0.05)$) of chat-tuned LLaMa-2 models when the persona is induces through the system message (as was done in the main text) compared to the user message input (as was done for other models).
-LLaMa-2 models do not exhibit value stability in either setting.
-This implies that the low stability is due to the LLaMa models themselves, and not due to the choice of a prompting template.
-((a)) Universalism
-((b)) Benevolence
-((c)) Power
-((d)) Achievement
-Figure 14: Rank-order stability between value expression (on the PVQ questionnaire) and the donation amount (on a downstream task).
-For more stable modes, donations are correlated with Universalism and Benevolence and negatively correlated with power and achievement.
+Appendix E Statistical analyses
+In the following section, we provide statistical analysis results for the experiments in the main text.
+In our experiments, we compare the stability of different models.
+We conduct the student’s t-test [60] on each pair of models with $p=0.05$.
+Given that we evaluate a total of 21 models, this amounts to a total of $\binom{21}{2}==210$ comparisons.
+We use the False Discovery Rate[61] to adjust the p-values to control for the number of comparisons.
+Figs 17, 16, 18, 19, 20, and 21 present the statistical analysis corresponding to experiments from Figs 4(b), 4(a), 5, 6, 8(b), 8(c), and 8(a), respectively.
+Figure 16: Statistical comparison of models’ Rank-order value stability for LLMs simulating fictional characters.
+This accompanies results shown in Fig 4(a).
+Black cells denote statistically significant difference between models.
+Figure 17: Statistical comparison of models’ Rank-order value stability for LLMs simulating real-world personas.
+This accompanies results shown in Fig 4(b).
+Black cells denote statistically significant difference between models.
+Figure 18: Statistical comparison of models’ Ipsative value stability for LLMs without the persona setting instructions.
+This accompanies results shown in Fig 5. Black cells denote statistically significant difference between models.
+Figure 19: Statistical comparison of models’ Rank-order value stability on the downstream Donation task.
+This accompanies results shown in Fig 8(b). Black cells denote statistically significant difference between models.
+Figure 20: Statistical comparison of models’ Rank-order value stability on the downstream Stealing task.
+This accompanies results shown in Fig 8(c). Black cells denote statistically significant difference between models.
+Figure 21: Statistical comparison of models’ Rank-order value stability on the downstream Religion task.
+This accompanies results shown in Fig 8(a). Black cells denote statistically significant difference between models.